### v1

# Chinese Academy of Sciences, Beijing 100190, China

Abstract
The primordial neutrino asymmetry leave profound imprints on the evolution history of the universe, which can be constrained by cosmological observations, including Big Bang Nucleosynthesis (BBN), Cosmic Microwave Background (CMB), and Large-Scale Structure (LSS). We present comprehensive analysis on implications and constraints of the primordial neutrino asymmetry $\xi_{\nu}$, based on a precise treatment of neutrino decoupling by solving the complete (anti)neutrino quantum kinetic equations in the Closed-Time-Path formalism. The effective number of neutrinos $N_{\rm eff}$ and (anti)neutrino spectral distortions are calculated, and we find that the non-instantaneous decoupling correction is given by $\delta N_{\rm eff}=0.0440+0.0102\,\xi_{\nu}^{2}$. Then we perform the state-of-the-art calculation for the abundance of light elements including (anti)neutrino spectral distortions, which indicate a positive asymmetry $0.032\leq\xi_{\nu}\leq 0.052$ from EMPRESS data. The implications of the neutrino asymmetry for the CMB and LSS are studied in detail, and we find that the Baryon Acoustic Oscillations (BAO) are also significantly affected by $\xi_{\nu}$ in addition to the sum of neutrino masses. A combined analysis with EMPRESS BBN, Planck CMB and BOSS BAO data yields a tighter constraint $\xi_{\nu}=0.024\pm 0.013$, which provides constraints on UV models capable of producing large asymmetries.
1 Introduction
[…]
A precise understanding of the neutrino decoupling process is essential for accurate constraints on neutrino properties. Notably, the partial reheating of neutrinos and antineutrinos due to the small overlap between neutrino decoupling and $e^{\pm}$ annihilations results in slightly non-thermal spectra and an increased total neutrino energy density, typically described by the effective number of relativistic species, $N_{\rm eff}$. The values of $N_{\rm eff}$ and the spectral distortions of neutrinos and antineutrinos are used as initial conditions to evaluate the impact of relic neutrinos in subsequent processes. For the Standard Model case, where asymmetries between neutrinos and antineutrinos are assumed to be negligible, significant progress has been made in detailing the neutrino decoupling process over recent decades Hannestad:1995rs ; Dolgov:1997mb ; Mangano:2001iu ; Mangano:2005cc ; Grohs:2015tfy ; deSalas:2016ztq ; Akita:2020szl ; Froustey:2019owm ; EscuderoAbenza:2020cmq ; Froustey:2020mcq ; Bennett:2020zkv . Including flavor oscillations, matter effects, finite-temperature QED corrections up to $\mathcal{O}(3)$, and full neutrino-electron and neutrino-neutrino collision terms, the current standard prediction yields $N_{\rm eff}(0)=3.0440\pm 0.0002$ Froustey:2020mcq ; Bennett:2020zkv .
On the other hand, inspired by the recent EMPRESS survey on primordial helium abundance empress and its potential to alleviate cosmological tensions Barenboim:2016lxv ; Yeung:2020zde ; Seto:2021tad ; Kumar:2022vee ; Yeung:2024krv , the scenario where asymmetries between neutrinos and antineutrinos are not negligible has recently attracted significant attention. Due to the sphaleron processes in the early universe Kuzmin:1985mm ; Khlebnikov:1988sr ; Harvey:1990qw ; Dreiner:1992vm , one might naively expect neutrino asymmetries, or lepton asymmetries, to be of the same order as the baryon asymmetry of the universe, which is strongly constrained by BBN and CMB observations: $\eta_{B}\equiv(n_{B}-n_{\bar{B}})/n_{\gamma}=(6.14\pm 0.04)\times 10^{-10}$ Planck:2018vyg , where $n_{B}$, $n_{\bar{B}}$, and $n_{\gamma}$ denote the number densities of baryons, antibaryons, and photons, respectively. However, substantial lepton asymmetries compared to the baryon asymmetry before the neutrino decoupling epoch can still be generated by various models, see $e.g.,$ Casas:1997gx ; Dolgov:1989us ; Bajc:1997ky ; Asaka:2005pn ; Asaka:2005an ; Pilaftsis:2003gt ; Borah:2022uos ; Kawasaki:2002hq ; Kawasaki:2022hvx .
The lepton asymmetries prior to the neutrino decoupling epoch, hereafter referred to as primordial neutrino asymmetries, have significant implications for the evolution of the universe. During the neutrino decoupling epoch, primordial neutrino asymmetries lead to an increased $N_{\rm eff}$ and specific spectral distortions of neutrinos and antineutrinos Grohs:2016cuu . Furthermore, if the asymmetries among different flavors are different, neutrinos and antineutrinos can undergo synchronous oscillations Bell:1998ds ; Pastor:2001iu ; Abazajian:2002qx ; Wong:2002fa , resulting in the redistribution of asymmetries among flavors Dolgov:2002ab ; Pastor:2008ti ; Mangano:2010ei ; Mangano:2011ip ; Castorina:2012md ; Barenboim:2016shh ; Johns:2016enc ; Froustey:2021azz . After neutrino decoupling, during the BBN epoch, primordial neutrino asymmetries primarily affect the final positron-to-neutron ratio and thus the final abundances of light elements. This effect mainly depends on the density distributions of electron neutrinos and antineutrinos, while corrections from the increased $N_{\rm eff}$ are subleading Sarkar:1995dd ; Iocco:2008va ; Pitrou:2018cgg ; Serpico:2005bc ; Mangano:2011ip ; Chu:2006ua ; Simha:2008mt ; Saviano:2013ktj ; Burns:2023sgx . Additionally, primordial neutrino asymmetries influence CMB anisotropies and LSS through increased $N_{\rm eff}$ and modified light element abundances, alongside the effects of massive neutrinos. Consequently, observations of BBN, CMB, and LSS can be utilized to constrain primordial neutrino asymmetries Simha:2008mt ; Oldengott:2017tzj ; Pitrou:2018cgg ; Matsumoto:2022tlr ; Burns:2022hkq ; Escudero:2022okz . In particular, recent studies from the EMPRESS survey indicate a preference for a positive asymmetry in electron neutrinos Matsumoto:2022tlr ; Burns:2022hkq ; Escudero:2022okz . Specifically, combining BBN and CMB observations yields $\xi_{\nu_{e}}=0.034\pm 0.014$ Escudero:2022okz , where $\xi_{\nu_{e}}=\mu_{\nu_{e}}/T_{\nu}$ represents the degeneracy parameter for electron neutrinos.
However, the current situation is less satisfactory compared to the Standard Model (SM) case due to persistent ambiguities regarding neutrino decoupling. These ambiguities particularly involve neutrino spectral distortions and the effects of flavor oscillations and QED corrections. As a result, the effects of primordial neutrino asymmetries on BBN and CMB, as well as the constraints derived from observational data, are generally evaluated using approximations. These approximations often include the instantaneous neutrino decoupling approximation and the assumption of massless neutrinos. To derive precise and accurate constraints on primordial neutrino asymmetries from both current and future observations, it is essential to rigorously analyze the implications of these asymmetries through a comprehensive treatment of the neutrino decoupling process. A recent study Froustey:2024mgf makes effort toward a rigorous analysis of primordial neutrino asymmetries. This study focuses on the implications of neutrino synchronous oscillations for BBN, while ignoring the correction from (anti)neutrino spectral distortions, etc, and furthermore, the implications for both the CMB and LSS are not addressed at all.
In this work, we focus on evaluating the neutrino decoupling process and its subsequent effects on BBN, CMB and LSS in the presence of primordial neutrino asymmetries. To accurately describe the evolution of neutrino density matrices during decoupling, including flavor oscillations and full collision effects, it is crucial to employ the complete set of neutrino quantum kinetic equations (QKEs). These equations can be derived using various approaches, such as perturbative expansions of the density matrix Sigl:1993ctk , relativistic generalizations of the Bogoliubov-Born-Green-Kirkwood-Yvon (BBGKY) equations Volpe:2013uxl ; Volpe:2015rla ; Froustey:2020mcq , and more generally from the Closed-Time Path (CTP) formalism of non-equilibrium quantum field theory Vlasenko:2013fja ; Blaschke:2016xxt . The main improvements in our study are as follows:
  • We re-derive the most general form of the QKEs for neutrinos and antineutrinos using the CTP formalism, incorporating primordial neutrino asymmetry.
  • We numerically solve the QKEs for the neutrino decoupling process in the presence of primordial neutrino asymmetry, including effects of flavor oscillations, matter interactions, finite-temperature QED (FTQED) corrections up to $\mathcal{O}(3)$ order, and complete neutrino-electron and neutrino-neutrino collision terms. This approach matches the accuracy used in current state-of-the-art predictions for the Standard Model (SM) case Froustey:2020mcq ; Bennett:2020zkv  [^1] Additional QED corrections, such as those to interaction vertices Cielo:2023bqp ; Jackson:2023zkl ; Drewes:2024wbw , are still debated for the SM case and are not included in this study..
  • We use the resulting density matrices to study the impacts of primordial neutrino asymmetry on the BBN, CMB, and LSS. The effects of neutrino spectral distortions are also highlighted by comparing these results with those obtained using the Fermi-Dirac distribution for neutrinos.
  • With the observation that the Baryon Acoustic Oscillations (BAO) are also affected by the primordial neutrino asymmetry, we derive the constraints on the primordial neutrino asymmetry by combining the traditional BBN and CMB datasets with the BAO observations for the first time.
Figure 1: A schematic illustration outlining the structure of this paper. Section 2 presents the derivation of the QKEs for neutrinos within the CTP formalism. These QKEs are subsequently applied to study neutrino decoupling in the presence of primordial neutrino asymmetry $\xi_{\nu}$ in Section 3. Based on such precise treatment of neutrino decoupling, Sections 4 and 5 explore the effects of primordial neutrino asymmetry on BBN, CMB and LSS. For each section, the primary input and output physical quantities that depend on $\xi_{\nu}$ are labelled in each box. Finally, in Section 6, cosmological constraints on $\xi_{\nu}$ are obtained by the MCMC analyses combining the theoretical predictions and observational data from BBN, CMB, and LSS.
The key physical results of this study are:
  • We numerically evaluate $N_{\rm eff}$ and the (anti)neutrino spectral distortions, incorporating primordial neutrino asymmetries with state-of-the-art accuracy. We find that $N_{\rm eff}$ can be expressed as $N_{\rm eff}=N_{\rm eff}^{\rm SM}+\delta N_{\rm eff}^{\xi_{\nu},\rm ther}+0.0102\,\xi_{\nu}^{2}$, where $N_{\rm eff}^{\rm SM}=3.0440\pm 0.002$ represents the SM prediction with comparable precision Froustey:2020mcq ; Bennett:2020zkv . The term $\delta N_{\rm eff}^{\xi_{\nu},\rm ther}$ accounts for the contribution from primordial neutrino asymmetry under the instantaneous decoupling approximation, and is given by $\delta N_{\rm eff}^{\xi_{\nu},\rm ther}\simeq 3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right)$, while the final term represents the additional correction from non-instantaneous decoupling effects in the presence with non-zero primordial neutrino asymmetry.
  • We provide state-of-the-art predictions for light element abundances produced by BBN, through which the EMPRESS measurement on helium abundance $Y_{P}|_{\rm EMPRESS}=0.2370^{+0.0034}_{-0.0033}$ immediately indicate a positive primordial neutrino asymmetry $0.032\leq\xi_{\nu}\leq 0.052$ with fixed baryon abundance $\omega_{b}$. We also find that the (anti)neutrino spectral distortions are important for precise predictions for BBN, especially for the abundance of Helium-4 $Y_{P}$.
  • For the constraints on primordial neutrino asymmetry, combing EMPRESS BBN and Planck CMB data yields $\xi_{\nu}=0.028\pm 0.013$, while incorporating BOSS BAO data in the analysis yields $\xi_{\nu}=0.024\pm 0.013$. Compared to previous result obtained in ref $\xi_{\nu}=0.034\pm 0.014$, our results indicate a $\sim 18\%$ and $\sim 29\%$ smaller central value and a slight reduction in the uncertainty, respectively.
  • The impacts of treating the sum of neutrino mass $\sum m_{\nu}$ and the primordial neutrino asymmetry $\xi_{\nu}$ as free parameters on the resulting constraints for each of them are also discussed. Finally, we show how the constraints on $\xi_{\nu}$ can be converted into the constraints on UV model parameters with a specific Q-ball decay model capable of producing large primordial neutrino asymmetries.
This paper is organized as follows. In Section 2, we re-derive the QKEs for Dirac and Majorana neutrinos using the CTP formalism. In Section 3, we present the QKEs in the context of the early universe and provides corresponding numerical results considering primordial neutrino asymmetries. The impacts of primordial neutrino asymmetries on the BBN epoch is discussed in Section 4, while their effects on both CMB and LSS are analyzed in Section 5. In Section 6, we present the results from a joint MCMC analysis of neutrino primordial asymmetries and neutrino masses, utilizing datasets from BBN, CMB and BAO observations. Finally, we conclude in Section 7. To guide the readers, a schematic illustration of this work are presented in Fig. 1. Throughout this manuscript, we use the natural units where $\hbar=c=k_{B}=1$.
2 Quantum Kinetic Equations for Dirac and Majorana Neutrinos
In this section, we derive the quantum kinetic equations (QKEs) for both Dirac and Majorana neutrinos using the closed-time-path (CTP) formalism of non-equilibrium quantum field theory. The QKEs generalize the traditional Boltzmann equation by governing the evolution of density matrices and incorporating both the phase-space distribution functions for the eigenstates and the quantum coherence between them. These equations have also been derived through various approaches and approximations in the literature (see $e.g.,$  Sigl:1993ctk ; Volpe:2013uxl ; Volpe:2015rla ; Froustey:2020mcq ).
2.1 The Closed-Time-Path formalism
The Closed-Time-Path (CTP) formalism, also known as the in-in formalism, provides a general framework for non-equilibrium quantum field theory (QFT). This approach is necessary because, in non-equilibrium situations, the final state of the system is not known a priori. The CTP formalism employs a closed time contour to handle such cases. Specifically, the path integral in this formalism is defined over a closed time path that starts at an initial time $t_{0}$, extends to infinity on the positive branch, and then returns to $t_{0}$ on the negative branch. This closed time path allows for the consistent treatment of quantum fields in non-equilibrium scenarios by incorporating both forward and backward evolution in time. Accordingly, depending on the location of the time arguments on the $\pm$ branches, there are four kinds of real-time Green functions, which we denote as
$\left(S_{ab}\right)^{>}(x,y)$ $\equiv\langle\nu_{a}(x_{-})\bar{\nu}_{b}(y_{+})\rangle~{},$ (1a)
$\left(S_{ab}\right)^{<}(x,y)$ $\equiv\langle\bar{\nu}_{b}(y_{-})\nu_{a}(x_{+})\rangle~{},$ (1b)
$\left(S_{ab}\right)^{\rm T}(x,y)$ $\equiv\langle\mathcal{T}\left(\nu_{a}(x_{+})\bar{\nu}_{b}(y_{+})\right)\rangle~{},$ (1c)
$\left(S_{ab}\right)^{\rm\bar{T}}(x,y)$ $\equiv\langle\bar{\mathcal{T}}\left(\nu_{a}(x_{-})\bar{\nu}_{b}(y_{-})\right)\rangle~{},$ (1d)
where $a$ and $b$ denote flavor indices of the neutrinos, while spinor indices are suppressed. The operators $\mathcal{T}$ and $\bar{\mathcal{T}}$ represent the usual time-ordering and reversed time-ordering operations, respectively. It is evident that only two of these four Green functions are independent. For example, by explicitly writing out the time-ordering operators, we obtain
$\left(S_{ab}\right)^{\rm T}(x,y)$ $=\theta(x^{0}-y^{0})\left(S_{ab}\right)^{>}(x,y)-\theta(y^{0}-x^{0})\left(S_{ab}\right)^{<}(x,y)~{},$ (2a)
$\left(S_{ab}\right)^{\rm\bar{T}}(x,y)$ $=\theta(y^{0}-x^{0})\left(S_{ab}\right)^{>}(x,y)-\theta(x^{0}-y^{0})\left(S_{ab}\right)^{<}(x,y)~{}.$ (2b)
In practice, it is also convenient to introduce the other two dependent propagators to describe the non-equilibrium systems of neutrinos, namely the statistical and spectral functions, defined as
$S^{F}_{ab}(x,y)$ $\equiv\frac{1}{2}\langle[\nu_{a}(x),\bar{\nu}_{b}(y)]\rangle=\frac{1}{2}\left(\left(S_{ab}\right)^{>}(x,y)-\left(S_{ab}\right)^{<}(x,y)\right)\,,$ (3a)
$S^{\mathcal{A}}_{ab}(x,y)$ $\equiv\frac{i}{2}\langle\{\nu_{a}(x),\bar{\nu}_{b}(y)\}\rangle=\frac{i}{2}\left(\left(S_{ab}\right)^{>}(x,y)+\left(S_{ab}\right)^{<}(x,y)\right)\,.$ (3b)
The advantage of this definition is that the statistical and spectral functions have straightforward physical interpretations Berges:2004yj ; Blaschke:2016xxt ; Vlasenko:2013fja . Specifically, the spectral function provides information about the spectrum of the theory, while the statistical function encodes information about the occupation numbers and quantum coherence of the available states, $i.e.,$ information about the density matrices. Additionally, we note that we do not specify whether the neutrinos are Dirac or Majorana particles here; the distinction between Dirac and Majorana neutrinos will be discussed later.
Therefore, the starting point for the QKEs is the equation of motion for the Green functions $S_{ab}(x,y)$, which can be expressed through the Dyson-Schwinger equation:
$\left(i\not{\partial}^{x}-m\right)\tilde{S}(x,y)=\mathbb{1}\,i\delta^{4}(x-y)+i\int d^{4}z\,\tilde{\Sigma}(x,z)\tilde{S}(z,y)\,.$ (4)
Here, $\tilde{S}$ denotes the matrix of Green functions in the Closed-Time-Path (CTP) space, and is defined as
$\tilde{S}=\begin{pmatrix}\left(S_{ab}\right)^{T}&\left(S_{ab}\right)^{<}\\ \left(S_{ab}\right)^{>}&-\left(S_{ab}\right)^{\bar{T}}\end{pmatrix}\,.$ (5)
In this context, $m\equiv\mathbb{1}_{2\times 2}\,m_{ab}$, where $m_{ab}$ is the neutrino mass matrix. The identity matrix preceding the Dirac delta function $\delta^{4}(x-y)$ is both an identity matrix in the CTP space and in the flavor space, i.e., $\mathbb{1}\,i\delta^{4}(x-y)=\mathbb{1}_{2n_{f}\times 2n_{f}}\,i\delta^{4}(x-y)$, with $n_{f}$ representing the number of neutrino flavors. Additionally, $\tilde{\Sigma}(x,y)$ is the neutrino proper self-energy, which is generally a functional of the Green functions and can be decomposed into a local singular term and a matrix in CTP space:
$\tilde{\Sigma}(x,y)=-\mathbb{1}_{2\times 2}i\Sigma(x)\delta^{4}(x-y)+\tilde{\Pi}\,,\quad\tilde{\Pi}=\begin{pmatrix}\left(\Pi_{ab}\right)^{T}&\left(\Pi_{ab}\right)^{<}\\ \left(\Pi_{ab}\right)^{>}&-\left(\Pi_{ab}\right)^{\bar{T}}\end{pmatrix}\,.$ (6)
In principle, $\Sigma(x)$ arises from the forward scattering potentials contributed by one-loop diagrams, while $\Pi^{\pm}(x,y)$ represents contributions from non-forward scatterings and annihilations that occur in at least two loops.
In the following, we suppress the flavor and spinor indices of the Green functions for simplicity. Expanding Eq. (4), the Dyson-Schwinger equation for $S^{\gtrless}$ can be written as
$\left(i\not{\partial}^{x}-m-\Sigma(x)\right)S^{\gtrless}(x,y)=i\int d^{4}z\left(\Pi^{H}(x,z)S^{\gtrless}(z,y)+\Pi^{\gtrless}(x,z)S^{H}(z,y)\right)\pm i\,\mathcal{C}_{\rm coll}(x,y)\,,$ (7)
which are known as the Kadanoff–Baym (KB) equations. Here, $S^{H}\equiv(S^{T}-S^{\bar{T}})/2$ and similarly for $\Pi^{H}$, while the collision term $\mathcal{C}_{\rm coll}$ is defined as
$\mathcal{C}_{\rm coll}(x,y)\equiv\frac{1}{2}\int d^{4}z\left(\Pi^{>}(x,z)S^{<}(z,y)+\Pi^{<}(x,z)S^{>}(z,y)\right)\,,$ (8)
[…]
Using the definition in Eq. (3), we can derive the equations for the statistical function $S^{F}$ and the spectral function $S^{\mathcal{A}}$:
$\left(i\not{\partial}^{x}-m-\Sigma(x)\right)S^{F}(x,y)$ $=i\int d^{4}z\left(\Pi^{H}(x,z)S^{F}(z,y)+\Pi^{F}(x,z)S^{H}(z,y)\right)-i\,\mathcal{C}_{\rm coll}(x,y)\,,$ (9a)
$\left(i\not{\partial}^{x}-m-\Sigma(x)\right)S^{\mathcal{A}}(x,y)$ $=i\int d^{4}z\left(\Pi^{H}(x,z)S^{\mathcal{A}}(z,y)+\Pi^{\mathcal{A}}(x,z)S^{H}(z,y)\right)\,.$ (9b)
We note that Eq. (9a) is equivalent to Eq. (26) for the statistical function given in Ref. Vlasenko:2013fja . As will be showed, our formulation is more convenient for the Wigner transform.
The purpose of the Wigner transform is to separate the evolution of the Green functions on the microscopic scale with respect to $r\equiv x-y$ from the evolution on the macroscopic scale with respect to $X\equiv(x+y)/2$, which is our primary focus. The (inverse) Wigner transform for any two-point (or self-energy) function $F(x,y)$ is given by
$F(k,X)$ $=\int d^{4}r\,e^{ik\cdot r}\,F\left(X+\frac{r}{2},X-\frac{r}{2}\right)\,,$ (10a)
$F(x,y)$ $=\int\frac{d^{4}k}{(2\pi)^{4}}\,e^{-ik\cdot(x-y)}\,F\left(k,\frac{x+y}{2}\right)\,.$ (10b)
To transform the equations for statistical and spectral functions, we also need the Wigner transform of the generalized convolution term like $\int d^{4}z\,F(x,z)\,G(z,y)$. The result is given by
$\int d^{4}z\,F(x,z)\,G(z,y)=\int\frac{d^{4}k}{(2\pi)^{4}}\,e^{-ik\cdot(x-y)}\,e^{-i\Diamond}\left\{F(k,X)\right\}\left\{G(k,X)\right\}\,,$ (11)
where the right-hand side is known as the Moyal product, and the diamond operator $\Diamond$ is defined as
$\Diamond\left\{F(k,X)\right\}\left\{G(k,X)\right\}=\frac{1}{2}\left(\frac{\partial F}{\partial X^{\mu}}\frac{\partial G}{\partial k_{\mu}}-\frac{\partial F}{\partial k_{\mu}}\frac{\partial G}{\partial X^{\mu}}\right)\,.$ (12)
The Wigner-transformed equations for the statistical and spectral functions, given by Eq. (9), can then be straightforwardly obtained as
[…]
$\quad+i\,e^{-i\Diamond}\left\{\Pi^{F}(k,X)\right\}\left\{S^{H}(k,X)\right\}-i\,\mathcal{C}_{\rm coll}(k,X)\,,$ (13b)
[…]
$\quad+i\,e^{-i\Diamond}\left\{\Pi^{\mathcal{A}}(k,X)\right\}\left\{S^{H}(k,X)\right\}\,,$
with the collision term
$\mathcal{C}_{\rm coll}(k,X)=\frac{1}{2}\left(e^{-i\Diamond}\left\{\Pi^{>}(k,X)\right\}\left\{S^{<}(k,X)\right\}-e^{-i\Diamond}\left\{\Pi^{<}(k,X)\right\}\left\{S^{>}(k,X)\right\}\right)\,.$ (14)
2.2 Gradient expansion of the kinetic equation
So far, the equations for statistical and spectral functions (13) involve infinite gradient expansions and are thus valid to all orders. In practice, we follow the procedure outlined in Ref. Blaschke:2016xxt and truncate these infinite series at the leading non-trivial order in a small parameter $\epsilon$. To achieve this, we first need to specify the power counting in the regime of interest, as follows:
  • The variation of physical quantities with respect to the macroscopic coordinate $X$ is small compared to its intrinsic de Broglie frequency, so each derivative $\partial_{X}$ carries one power of $\epsilon$.
  • Compared to the neutrino energy, the masses and interaction potentials are small, so terms involving $m$ or $\Sigma$ carry one power of $\epsilon$.
  • Contributions to the self-energy $\Pi$ appear at least at the 2-loop order in the Feynman diagram expansion, so terms involving $\Pi$ carry two power of $\epsilon$.
  • The variation of the two-point functions from equilibrium is small, so $S^{H}=(S^{T}-S^{\bar{T}})/2$ carries one power of $\epsilon$.
In conclusion, our power counting reads
$\frac{\partial_{X},m,\Sigma,S^{H}}{E_{\nu}}=O(\epsilon)\,,\quad\frac{\tilde{\Pi}^{\pm}}{E_{\nu}}=O(\epsilon^{2})\,,$ (15)
where $E_{\nu}$ is the neutrino energy. Additionally, we note that the contribution from the $\Pi^{H}$ term can be absorbed into the $\Sigma$ term at the considered order, so we will omit the relevant terms in the following equations.
Next, we expand the equations for statistical and spectral functions (13) order by order to derive the QKEs for neutrinos and antineutrinos. To do this, we first note that the kinetic equation (13b) involves the sixteen spinor components (scalar, pseudoscalar, vector, axial-vector, tensor) of the statistical function and the self-energies, which can be decomposed as follows:
$S^{F}$ $=\left[F_{S}+\left(F_{V}^{R}\right)^{\mu}\gamma_{\mu}-\frac{i}{4}\left(F_{T}^{L}\right)^{\mu\nu}\sigma_{\mu\nu}\right]P_{L}+\left[F_{S}^{\dagger}+\left(F_{V}^{L}\right)^{\mu}\gamma_{\mu}+\frac{i}{4}\left(F_{T}^{R}\right)^{\mu\nu}\sigma_{\mu\nu}\right]P_{R}\,,$ (16a)
$\Pi^{F}$ $=\left[\Pi_{S}+\Pi_{R}^{\mu}\gamma_{\mu}-\frac{i}{4}\left(\Pi_{T}^{L}\right)^{\mu\nu}\sigma_{\mu\nu}\right]P_{L}+\left[\Pi_{S}^{\dagger}+\Pi_{L}^{\mu}\gamma_{\mu}+\frac{i}{4}\left(\Pi_{T}^{R}\right)^{\mu\nu}\sigma_{\mu\nu}\right]P_{R}\,,$ (16b)
where $P_{L,R}\equiv\frac{1\mp\gamma_{5}}{2}$ and $\sigma_{\mu\nu}\equiv\frac{i}{2}[\gamma_{\mu},\gamma_{\nu}]$. Due to the hermiticity conditions, we have $F_{V}^{L\dagger}=F_{V}^{L}$, $F_{V}^{R\dagger}=F_{V}^{R}$, and $F_{T}^{L\dagger}=F_{T}^{R}$, with similar conditions applying to the components of $\Pi^{F}$. An analogous decomposition for the forward scattering potential $\Sigma(x)$ is also permissible.
Therefore, the kinetic equation (13b) encompasses both the QKEs and the algebraic constraints for various spinor components. The basic strategy to obtain the QKEs is as follows:
  1. 1.
    Solve Eq. (13) at $\mathcal{O}(\epsilon^{0})$ to obtain the leading order solutions for the statistical and spectral functions, and establish the relations between the remaining spinor components of the statistical function and the density matrices for Dirac and Majorana neutrinos.
  2. 2.
    Beyond $\mathcal{O}(\epsilon^{0})$, all spinor components of $S^{F}$ receive additional corrections due to neutrino masses and interactions. Thus, solve the kinetic equation (13b) at $\mathcal{O}(\epsilon^{1})$ to find the constraint relations among these small components.
  3. 3.
    Expand the kinetic equation (13b) to $\mathcal{O}(\epsilon^{2})$ and reorganize it using the constraint relations obtained at $\mathcal{O}(\epsilon^{1})$.
  4. 4.
    Finally, extract the QKEs for Dirac and Majorana neutrinos, incorporating the relations between the spinor components of the statistical function and the density matrices.
The detailed derivation of these steps is presented in the following sub-subsections.
2.2.1 Kinetic equation to $\mathcal{O}(\epsilon^{0})$
[…]
$\left(F_{V}^{L,R}\right)^{\mu}(k,x)$ $=\hat{\kappa}^{\mu}(k)\,F_{L,R}(k,x)\,,$ (17a)
$\left(F_{T}^{L}\right)_{\mu\nu}(k,x)$ $=e^{-i\varphi(k)}(\hat{\kappa}(k)\wedge\hat{x}^{-}(k))_{\mu\nu}\,\Phi(k,x)\,,$ (17b)
$\left(F_{T}^{R}\right)_{\mu\nu}(k,x)$ $=e^{i\varphi(k)}(\hat{\kappa}(k)\wedge\hat{x}^{+}(k))_{\mu\nu}\,\Phi^{\dagger}(k,x)\,.$ (17c)
Here, we have introduced a set of basis vectors used to express the Lorentz tensors and components of the two-point functions. The basis consists of two light-like vectors $\hat{\kappa}^{\mu}(k)=({\rm sign}(k^{0}),\hat{k})$ and $\hat{\kappa}^{\prime\mu}(k)=({\rm sign}(k^{0}),-\hat{k})$, and two transverse four-vectors $\hat{x}_{1,2}(k)$ which satisfy the relations $\hat{\kappa}\cdot\hat{\kappa}=\hat{\kappa}^{\prime}\cdot\hat{\kappa}^{\prime}=0$, $\hat{\kappa}\cdot\hat{\kappa}^{\prime}=2$, $\hat{\kappa}\cdot\hat{x}_{i}=\hat{\kappa}^{\prime}\cdot\hat{x}_{i}=0$, and $\hat{x}_{i}\cdot\hat{x}_{j}=-\delta_{ij}$. Additionally, we define $\hat{x}^{\pm}\equiv\hat{x}_{1}\pm i\hat{x}_{2}$, such that $\hat{x}^{+}\cdot\hat{x}^{-}=-2$.
In addition, $F_{L,R}(k,x)$ and $\Phi(k,x)$ can be organized into a $2n_{f}\times 2n_{f}$ matrix:
$\hat{F}=\left(\begin{array}[]{cc}F_{L}&\Phi\\ \Phi^{\dagger}&F_{R}\end{array}\right)~{}.$ (18)
The advantage of this organization is that it can be directly related to the conventional density matrices. At $\mathcal{O}(\epsilon^{0})$, the density matrices for neutrinos and antineutrinos can be obtained via integrals over positive and negative frequencies. For Dirac neutrinos, we have
$-2\int_{0}^{\infty}\frac{dk^{0}}{2\pi}\hat{F}(k,x)$ $=F(\vec{k},x)-\frac{1}{2}\mathbb{1}\,,$ (19a)
$-2\int_{-\infty}^{0}\frac{dk^{0}}{2\pi}\hat{F}(k,x)$ $=\bar{F}(-\vec{k},x)-\frac{1}{2}\mathbb{1}\,.$ (19b)
The density matrices $F$ and $\bar{F}$ are also $2n_{f}\times 2n_{f}$ matrices:
$F(\vec{p},x)=\left(\begin{array}[]{cc}\varrho_{LL}&\varrho_{LR}\\ \varrho_{RL}&\varrho_{RR}\end{array}\right)\,;\qquad\qquad\bar{F}(\vec{p},x)=\left(\begin{array}[]{cc}\bar{\varrho}_{RR}&\bar{\varrho}_{RL}\\ \bar{\varrho}_{LR}&\bar{\varrho}_{LL}\end{array}\right)\,,$ (20)
where $n_{f}$ is the total number of neutrino flavors, and each block is a square $n_{f}\times n_{f}$ matrix with $\varrho_{hh^{\prime}}=\varrho_{hh^{\prime}}^{ij}$, where $i,\,j$ denote the flavors and $h,\,h^{\prime}\in\{L,R\}$ denote the chirality. Among the various components, $\varrho_{LL}$ and $\bar{\varrho}_{RR}$ represent the density matrices for active left-handed neutrinos and right-handed antineutrinos, respectively. Conversely, $\varrho_{RR}$ and $\bar{\varrho}_{LL}$ correspond to sterile right-handed neutrinos and left-handed antineutrinos. The remaining density matrices account for the quantum coherence between different chiralities of neutrinos and antineutrinos across all flavors. Note that there is no single-particle density matrix for quantum coherence between Dirac neutrinos and antineutrinos, as they are fundamentally distinct particles.
For Majorana neutrinos, the antiparticle is the neutrino itself, and the density matrices for neutrinos and antineutrinos are related by $\varrho_{hh^{\prime}}^{ij}=\bar{\varrho}_{hh^{\prime}}^{ji}$ through a phase change. Therefore, only the density matrices for active neutrinos are necessary, which can be consolidated into a single $2n_{f}\times 2n_{f}$ matrix using the definitions $\varrho\equiv\varrho_{LL}$, $\bar{\varrho}\equiv\bar{\varrho}_{RR}=\varrho_{RR}^{T}$, and $\phi\equiv\varrho_{LR}$:
${\cal F}=\left(\begin{array}[]{cc}\varrho&\phi\\ \phi^{\dagger}&\bar{\varrho}^{T}\end{array}\right)~{},$ (21)
where $\varrho$ and $\bar{\varrho}$ denote the density matrices for active neutrinos and antineutrinos, respectively, and $\phi$ describes the quantum spin coherence between neutrinos and antineutrinos. Similar to the Dirac case, the density matrix for Majorana neutrinos, ${\cal F}$, can be extracted from the two-point functions via
$-2\int_{0}^{\infty}\frac{dk^{0}}{2\pi}\hat{F}(k,x)=\mathcal{F}(\vec{k},x)-\frac{1}{2}\mathbb{1}\,.$ (22)
Beyond $\mathcal{O}(\epsilon^{0})$, the equations (19) and (22) serve as the definitions of density matrices for Dirac and Majorana neutrinos, respectively, and are used to extract the QKEs from the evolution equation of $\hat{F}$. Additionally, to obtain the QKEs at $\mathcal{O}(\epsilon^{2})$, only the $\mathcal{O}(\epsilon^{0})$ expression for the vector component of the spectral function is required, which is given by $\hat{\rho}(k)=2i\pi|\vec{k}|\delta(k^{2})\text{sgn}(k^{0})$.
2.2.2 Kinetic equation to $\mathcal{O}(\epsilon^{1})$
Beyond $\mathcal{O}(\epsilon^{0})$, we still consider $F_{L,R}(k,x)$ and $\Phi(k,x)$ as the four independent spinor components of $S^{F}(k,x)$, which can be isolated using the projections:
$F_{L,R}(k,x)$ $\equiv\frac{1}{4}\mathrm{Tr}\left(\gamma_{\mu}P_{L,R}F^{(\nu)}(k,x)\right)\hat{\kappa}^{\prime\mu}(k)\,,$ (23a)
$\Phi^{(\dagger)}(k,x)$ $\equiv\mp\frac{i}{16}\mathrm{Tr}\left(\sigma_{\mu\nu}P_{L/R}F^{(\nu)}(k,x)\right)(\hat{\kappa}^{\prime}(k)\wedge\hat{x}^{\pm}(k))^{\mu\nu}e^{\pm i\varphi(k)}\,,$ (23b)
where the upper (lower) signs and indices refer to $\Phi$ ($\Phi^{\dagger}$). The dispersion relations for the functions $F_{L,R}(k,x)$ and $\Phi(k,x)$, as well as for the other spinor components of $S^{F}(k,x)$, are modified by neutrino masses and interactions at $\mathcal{O}(\epsilon^{1})$. Therefore, the general statistical function $S^{F}(k,x)$ can be expressed as
$S^{F}$ $=\left[F_{R}\hat{\kappa}^{\mu}\gamma_{\mu}-\frac{i}{4}e^{-i\varphi(k)}(\hat{\kappa}(k)\wedge\hat{x}^{-}(k))_{\mu\nu}\Phi\sigma_{\mu\nu}\right]P_{L}$ (24)
$+\left[F_{L}\hat{\kappa}^{\mu}\gamma_{\mu}+\frac{i}{4}e^{i\varphi(k)}(\hat{\kappa}(k)\wedge\hat{x}^{+}(k))_{\mu\nu}\Phi^{\dagger}\sigma_{\mu\nu}\right]P_{R}$
$+\left[\Delta_{S}+(\Delta_{R})^{\mu}\gamma_{\mu}-\frac{i}{4}(\Delta_{T}^{L})^{\mu\nu}\sigma_{\mu\nu}\right]P_{L}$
$+\left[\Delta_{S}^{\dagger}+(\Delta_{L})^{\mu}\gamma_{\mu}+\frac{i}{4}(\Delta_{T}^{R})^{\mu\nu}\sigma_{\mu\nu}\right]P_{R}\,,$
where $\Delta$ represents the $\mathcal{O}(\epsilon)$ small components.
Expand the kinetic equation for $S^{F}(k,x)$ to $O(\epsilon)$, we obtain
[…]
Substituting the general statistical function $S^{F}(k,x)$, including the $O(\epsilon)$ corrections, into the kinetic equation and decomposing into scalar, vector, and tensor components, we derive the dispersion relations for $F_{L,R}(k,x)$ and $\Phi(k,x)$ at $O(\epsilon)$:
$\left(k\cdot\hat{\kappa}\right)F_{R}-\frac{1}{2}\left\{\Sigma_{L}^{\kappa},F_{R}\right\}$ $=O\left(\epsilon^{2}\right),$ (26)
$\left(k\cdot\hat{\kappa}\right)F_{L}-\frac{1}{2}\left\{\Sigma_{R}^{\kappa},F_{L}\right\}$ $=O\left(\epsilon^{2}\right),$ (27)
$\left(k\cdot\hat{\kappa}\right)P_{+}^{ij}F_{T}^{j}-\frac{1}{2}\left(\Sigma_{R}^{\kappa}P_{+}^{ij}F_{T}^{j}+P_{+}^{ij}F_{T}^{j}\Sigma_{L}^{\kappa}\right)$ $=O\left(\epsilon^{2}\right),$ (28)
$\left(k\cdot\hat{\kappa}\right)P_{-}^{ij}F_{T}^{j}-\frac{1}{2}\left(\Sigma_{L}^{\kappa}P_{-}^{ij}F_{T}^{j}+P_{-}^{ij}F_{T}^{j}\Sigma_{R}^{\kappa}\right)$ $=O\left(\epsilon^{2}\right),$ (29)
and the expressions for the small components
$\Delta_{L/R}^{\kappa}$ $=O\left(\epsilon^{2}\right)\ \ \ \ \,,\ \ \ \ \ \Delta_{T}^{i}=O\left(\epsilon^{2}\right),$ (30)
$\Delta_{S}$ $=\frac{1}{2\left|\vec{k}\right|}\left(m^{\dagger}F_{R}+F_{L}m^{\dagger}\right)+\frac{P_{+}^{ij}}{2\left|\vec{k}\right|}\left(i\partial^{i}F_{T}^{j}-\left(\Sigma_{R}^{i}F_{T}^{j}-F_{T}^{j}\Sigma_{L}^{i}\right)\right)$ (31)
$\Delta_{T}$ $=-\frac{1}{2\left|\vec{k}\right|}\left(m^{\dagger}F_{R}-F_{L}m^{\dagger}\right)+\frac{P_{+}^{ij}}{2\left|\vec{k}\right|}\left(\Sigma_{R}^{i}F_{T}^{j}+F_{T}^{j}\Sigma_{L}^{i}\right),$ (32)
$\Delta_{L}^{i}$ $=\frac{1}{2\left|\vec{k}\right|}\left(m^{\dagger}P_{-}^{ij}F_{T}^{j}+P_{+}^{ij}F_{T}^{j}m\right)+\frac{1}{\left|\vec{k}\right|}\left(\frac{1}{2}\epsilon^{ij}\partial^{j}F_{L}-\left(P_{-}^{ij}\Sigma_{R}^{j}F_{L}+F_{L}P_{+}^{ij}\Sigma_{R}^{j}\right)\right),$ (33)
$\Delta_{R}^{i}$ $=-\frac{1}{2\left|\vec{k}\right|}\left(mP_{+}^{ij}F_{T}^{j}+P_{-}^{ij}F_{T}^{j}m^{\dagger}\right)-\frac{1}{\left|\vec{k}\right|}\left(\frac{1}{2}\epsilon^{ij}\partial^{j}F_{R}+\left(P_{+}^{ij}\Sigma_{L}^{j}F_{R}+F_{R}P_{-}^{ij}\Sigma_{L}^{j}\right)\right).$ (34)
2.2.3 Kinetic equation to $\mathcal{O}(\epsilon^{2})$
At $O(\epsilon^{2})$, the kinetic equation $S^{F}(k,x)$ becomes
[…]
Inserting the expressions for the small components obtained at $\mathcal{O}(\epsilon)$ and reorganizing the kinetic equations, we can derive the kinetic equation and the dispersion relation for $\hat{F}$ as
$\partial^{\kappa}\hat{F}+\frac{1}{2|\vec{k}|}\left\{\Sigma^{i},\partial^{i}\hat{F}\right\}+\frac{1}{2}\left\{\frac{\partial\Sigma^{\kappa}}{\partial x^{\mu}},\frac{\partial\hat{F}}{\partial k_{\mu}}\right\}$ $=-i\left[H,\hat{F}\right]+\hat{C}~{},$ (36a)
$\left\{\hat{\kappa}(k)\cdot k-\Sigma^{\kappa}\ ,\ \hat{F}\right\}$ $=0~{},$ (36b)
where we have defined $\partial^{\kappa}\equiv\hat{\kappa}(k)\cdot\partial$ and $\partial^{i}\equiv\hat{x}^{i}(k)\cdot\partial$, and $H$ is the Hamiltonian-like operator given by
$H=\left(\begin{array}[]{cc}H_{R}&H_{LR}\\ H_{LR}^{\dagger}&H_{L}\end{array}\right)~{},\qquad$ (37)
with
$H_{R}$ $=\Sigma_{R}^{\kappa}+\frac{1}{2|\vec{k}|}\left(m^{\dagger}m-\epsilon^{ij}\partial^{i}\Sigma_{R}^{j}+4\Sigma_{R}^{+}\Sigma_{R}^{-}\right)\,,$ (38a)
$H_{L}$ $=\Sigma_{L}^{\kappa}+\frac{1}{2|\vec{k}|}\left(mm^{\dagger}+\epsilon^{ij}\partial^{i}\Sigma_{L}^{j}+4\Sigma_{L}^{-}\Sigma_{L}^{+}\right)\,,$ (38b)
$H_{LR}$ $=-\frac{1}{|\vec{k}|}\left(\Sigma_{R}^{+}\,m^{\dagger}-m^{\dagger}\,\Sigma_{L}^{-}\right)~{},$ (38c)
where $\Sigma_{L,R}^{\pm}\equiv(1/2)\,e^{\pm i\varphi}\,(x_{1}\pm ix_{2})_{\mu}\,\Sigma^{\mu}_{L,R}$ and $\epsilon^{ij}$ is the two-dimensional Levi-Civita symbol. Finally, the collision term $\hat{C}$ is given by
$\hat{C}=-\frac{1}{2}\left\{\hat{\Pi}^{>},{\hat{S}}^{<}\right\}+\frac{1}{2}\left\{\hat{\Pi}^{<},{\hat{S}}^{>}\right\}\,,\$ (39)
with
[…]
and the $2n_{f}\times 2n_{f}$ self energy matrices $\hat{\Pi}^{\pm}(k)$ are given by
$\hat{\Pi}^{\pm}(k,x)$ $=\left(\begin{array}[]{cc}\Pi_{R}^{\kappa\pm}&2P_{T}^{\pm}\\ 2P_{T}^{\pm\dagger}&\Pi_{L}^{\kappa\pm}\end{array}\right)~{}.$ (43)
where the components are obtained by the projections
$\Pi_{L,R}^{\kappa}(k,x)$ $=\frac{1}{2}\hat{\kappa}_{\mu}\,\mathrm{Tr}\left[\tilde{\Pi}(k,x)\,\gamma^{\mu}P_{L,R}\right]\,,$ (44a)
$P_{T}(k,x)$ $=\frac{ie^{i\varphi}}{16}\,(\hat{\kappa}\wedge\hat{x}^{+})^{\mu\nu}\,\,\mathrm{Tr}\left[\tilde{\Pi}(k,x)\,\sigma_{\mu\nu}P_{R}\right]~{}.$ (44b)
2.3 QKEs for Dirac and Majorana neutrinos
Finally, integrating Eq. (36a) over positive and negative frequencies using Eq. (19), we obtain the quantum kinetic equations (QKEs) for Dirac neutrinos:
$\partial^{\kappa}F+\frac{1}{2|\vec{k}|}\left\{\Sigma^{i},\partial^{i}F\right\}-\frac{1}{2}\left\{\frac{\partial\Sigma^{\kappa}}{\partial\vec{x}},\frac{\partial F}{\partial\vec{k}}\right\}$ $=-i[H,F]+{\cal C}\,,$ (45a)
$\partial^{\kappa}\bar{F}-\frac{1}{2|\vec{k}|}\left\{\Sigma^{i},\partial^{i}\bar{F}\right\}+\frac{1}{2}\left\{\frac{\partial\Sigma^{\kappa}}{\partial\vec{x}},\frac{\partial\bar{F}}{\partial\vec{k}}\right\}$ $=-i[\bar{H},\bar{F}]+\bar{\cal C}\,.$ (45b)
The terms on the left-hand side (LHS) of the QKEs generalize the classical space-time derivative, drift, and force terms. The first term on the right-hand side (RHS) of the QKEs describes neutrino flavor oscillations, which are induced by neutrino masses and forward scattering between neutrinos and background particles. This corresponds to the generalization of the well-known Mikheyev-Smirnov-Wolfenstein (MSW) effect. The second term on the RHS represents the collision term with a non-trivial matrix structure:
${\cal C}$ $=\frac{1}{2}\left\{\Pi^{>},F\right\}-\frac{1}{2}\left\{\Pi^{<},\mathbb{1}-F\right\}\,,$ (46a)
$\bar{\cal C}$ $=\frac{1}{2}\left\{\bar{\Pi}^{>},\bar{F}\right\}-\frac{1}{2}\left\{\bar{\Pi}^{<},\mathbb{1}-\bar{F}\right\}\,,$ (46b)
where
[…]
On the other hand, the QKEs for Majorana neutrinos can be derived by integrating Eq. (36a) over positive frequencies using Eq. (22), yielding:
$\partial^{\kappa}{\cal F}+\frac{1}{2|\vec{k}|}\left\{\Sigma^{i},\partial^{i}{\cal F}\right\}-\frac{1}{2}\left\{\frac{\partial\Sigma^{\kappa}}{\partial\vec{x}},\frac{\partial{\cal F}}{\partial\vec{k}}\right\}$ $=-i[H,{\cal F}]+{\cal C}_{M}\,,$ (49)
with
${\cal C}_{M}$ $=\frac{1}{2}\left\{\Pi^{>},{\cal F}\right\}-\frac{1}{2}\left\{\Pi^{<},\mathbb{1}-{\cal F}\right\}\,.$ (50)
Explicitly, the potentials induced by forward scattering $\Sigma_{R,L}$ and the collision terms for Dirac and Majorana neutrinos should be calculated from the 1-loop and 2-loop self-energy diagrams, depending on the interaction terms for the neutrinos. For active neutrinos in the early universe, we will present the resulting forward potentials and collision terms in the next section.
3 Neutrino Decoupling in the Presence of Primordial Neutrino Asymmetry
3.1 QKEs in the early universe
In the early universe, both the Standard Model (SM) thermal bath and the neutrino system can be approximated as homogeneous and isotropic. Under these conditions, the QKEs for neutrinos can be significantly simplified. First, the drift and force terms on the left-hand side are eliminated, and a Hubble term is introduced to account for the expansion of the Universe. Second, the density matrix for neutrinos is considered to depend solely on the absolute value of the momentum. Finally, the coherence matrices between left-handed and right-handed neutrinos, namely $\varrho_{LR,RL}$ and $\bar{\varrho}_{LR,RL}$ for Dirac neutrinos, and $\phi$ for Majorana neutrinos, are reduced to zero, leading to corresponding collision terms also being zero. Consequently, the non-zero density matrices are $\varrho_{LL,RR}$ and $\bar{\varrho}_{LL,RR}$ for Dirac neutrinos, and $\varrho$ and $\bar{\varrho}$ for Majorana neutrinos.
For active neutrinos, the QKEs for $\varrho_{LL}$ and $\bar{\varrho}_{RR}$ in the case of Dirac neutrinos, are identical with those for $\varrho$ and $\bar{\varrho}$ for Majorana neutrinos. Therefore, in the subsequent analysis, we will not distinguish between Dirac and Majorana neutrinos and will use $\varrho$ and $\bar{\varrho}$ to represent the density matrices for neutrinos and antineutrinos, respectively. In the flavor basis, the QKEs can be written as:
$i\left[\frac{\partial}{\partial t}-Hp\frac{\partial}{\partial p}\right]\varrho$ $=\Big{[}H_{\nu},\varrho\Big{]}+i\mathcal{I}\,,$ (51)
$i\left[\frac{\partial}{\partial t}-Hp\frac{\partial}{\partial p}\right]\bar{\varrho}$ $=-\Big{[}\bar{H}_{\nu},\bar{\varrho}\Big{]}+i\bar{\mathcal{I}}\,,$ (52)
where $H\equiv\dot{a}/a$ is the Hubble rate, and $H_{\nu}$ and $\bar{H}_{\nu}$ are the Hamiltonian terms for neutrinos and antineutrinos, defined as:
$H_{\nu}$ $=U\frac{\mathbb{M}^{2}}{2p}U^{\dagger}+\sqrt{2}G_{F}\left(\mathbb{N}_{l}+\mathbb{N}_{\nu}\right)-2\sqrt{2}G_{F}p\left(\frac{\mathbb{E}_{l}+\mathbb{P}_{l}}{m_{W}^{2}}+\frac{4}{3}\frac{\mathbb{E}_{\nu}}{m_{Z}^{2}}\right)\,,$ (53)
$\bar{H}_{\nu}$ $=U\frac{\mathbb{M}^{2}}{2p}U^{\dagger}-\sqrt{2}G_{F}\left(\mathbb{N}_{l}+\mathbb{N}_{\nu}\right)-2\sqrt{2}G_{F}p\left(\frac{\mathbb{E}_{l}+\mathbb{P}_{l}}{m_{W}^{2}}+\frac{4}{3}\frac{\mathbb{E}_{\nu}}{m_{Z}^{2}}\right)\,.$ (54)
Here, $\mathbb{M}$ represents the neutrino mass matrix in the mass basis, and $U$ is the PMNS transformation matrix that relates the flavor eigenstates to the mass eigenstates. Using the standard parametrization of the PMNS matrix Giunti:2007ry ; Gariazzo:2019gyi , we have:
$\mathbb{M}=\begin{pmatrix}m_{1}^{2}&0&0\\ 0&m_{2}^{2}&0\\ 0&0&m_{3}^{2}\end{pmatrix}\,,\quad U=\begin{pmatrix}c_{12}c_{13}&s_{12}c_{13}&s_{13}\\ -s_{12}c_{23}-c_{12}s_{23}s_{13}&c_{12}c_{23}-s_{12}s_{23}s_{13}&s_{23}c_{13}\\ s_{12}s_{23}-c_{12}c_{23}s_{13}&-c_{12}s_{23}-s_{12}c_{23}s_{13}&c_{23}c_{13}\end{pmatrix}\,,$ (55)
where $c_{ij}=\cos{\theta_{ij}}$, $s_{ij}=\sin{\theta_{ij}}$, and $\theta_{ij}$ are the mixing angles. We neglect the potential CP-violating phase in this work.
The remaining terms in the Hamiltonian account for the matter potentials induced by coherent scatterings between neutrinos and other plasma particles. Here, $m_{W}$ and $m_{Z}$ denote the masses of the $W$ and $Z$ bosons, respectively, and $G_{F}$ is the Fermi constant. The second term in the Hamiltonian reflects the net number densities for charged leptons ($\mathbb{N}_{l}\equiv\mathrm{diag}(n_{e}-n_{\bar{e}},n_{\mu}-n_{\bar{\mu}},0)$) and neutrinos ($\mathbb{N}_{\nu}\equiv n_{\nu}-n_{\bar{\nu}}$). The third term accounts for the energy densities of charged leptons and neutrinos ($\mathbb{E}_{l}\equiv\mathrm{diag}(\rho_{e}+\rho_{\bar{e}},\rho_{\mu}+\rho_{\bar{\mu}},0)$, $\mathbb{P}_{l}\equiv\mathrm{diag}(P_{e}+P_{\bar{e}},P_{\mu}+P_{\bar{\mu}},0)$, $\mathbb{E}_{\nu}\equiv\rho_{\nu}+\rho_{\bar{\nu}}$). In practice, the net number densities for charged leptons are often negligible due to the small baryon asymmetry. However, the net number densities for neutrinos can lead to significant effects, especially if primordial neutrino asymmetries are present. This can result in pronounced neutrino flavor oscillations, including the well-known "synchronized oscillation" when this term dominates the Hamiltonian Bell:1998ds ; Pastor:2001iu ; Abazajian:2002qx ; Wong:2002fa ; Johns:2016enc ; Froustey:2021azz ; Froustey:2024mgf .
Finally, the collision terms $\mathcal{I}$ and $\bar{\mathcal{I}}$ account for contributions from neutrino-electron/positron scattering ($\nu e\leftrightarrow\nu e$), annihilation ($\nu\bar{\nu}\leftrightarrow e^{+}e^{-}$), and neutrino-neutrino self-interactions ($\nu\nu\leftrightarrow\nu\nu$, $\nu\bar{\nu}\leftrightarrow\nu\bar{\nu}$). These terms can be expressed as
$\mathcal{I}\equiv\mathcal{I}_{\rm sc}+\mathcal{I}_{\rm ann}+\mathcal{I}_{\nu\nu}$ (56)
with a similar decomposition for the collision term of antineutrinos $\bar{\mathcal{I}}$. The collision terms are generally functions of the neutrino density matrices, and their explicit forms are detailed in Appendix A. Contributions from other interaction processes, such as $\mu^{\pm}$ scattering and annihilation, are neglected due to their minimal impact at the temperature of neutrino decoupling.
In practice, comoving coordinates and corresponding comoving variables are employed to eliminate the Hubble expansion term in the QKEs. The comoving temperature is defined as $T_{\rm CM}\propto a^{-1}$ Grohs:2015tfy , with $T_{\rm CM}=T_{\nu}=T_{\gamma}$ at the initial time before neutrino decoupling when all species are tightly coupled. Using the comoving temperature, the reduced scale factor, the comoving momentum, and the dimensionless photon temperature are expressed as Mangano:2005cc ; Froustey:2020mcq ; Bennett:2020zkv
$x=m_{e}/T_{\rm CM}\,,\qquad y=p/T_{\rm CM}\,,\qquad z=T_{\gamma}/T_{\rm CM}\,.$ (57)
By definition, the dimensionless photon temperature $z$ equals 1 at high temperatures and increases beyond 1 later due to the annihilation of electrons and positrons into photons. For the convenience of the QKEs, we also define the dimensionless energy density and pressure as $\tilde{\rho}\equiv(x/m_{e})^{4}\rho$ and $\tilde{P}\equiv(x/m_{e})^{4}P$, respectively.
The comoving QKEs for neutrinos and antineutrinos then take the form as:
$\frac{{\rm d}\varrho(x,y)}{{\rm d}x}$ $=$ $\sqrt{\frac{3m^{2}_{\rm Pl}}{8\pi\tilde{\rho}_{\rm tot}}}\left\{-i\frac{x^{2}}{m_{e}^{3}}\left[\tilde{H}_{\nu},\varrho\right]+\frac{m_{e}^{3}}{x^{4}}\tilde{\mathcal{I}}(\varrho,\bar{\varrho})\right\}\,,$
$\frac{{\rm d}\bar{\varrho}(x,y)}{{\rm d}x}$ $=$ $\sqrt{\frac{3m^{2}_{\rm Pl}}{8\pi\rho}}\left\{-i\frac{x^{2}}{m_{e}^{3}}\left[\tilde{H}_{\bar{\nu}},\varrho\right]+\frac{m_{e}^{3}}{x^{4}}\tilde{\bar{\mathcal{I}}}(\bar{\varrho},\varrho)\right\}\,,$
with the comoving Hamiltonian terms
$\tilde{H}_{\nu}$ $=U\frac{\mathbb{M}^{2}}{2y}U^{\dagger}+\frac{\sqrt{2}G_{\rm F}m_{e}^{4}}{x^{4}}\,\mathbb{N}_{\nu}-\frac{2\sqrt{2}G_{\rm F}ym_{e}^{6}}{x^{6}}\left(\frac{\mathbb{E}_{\ell}+\mathbb{P}_{\ell}}{m_{W}^{2}}+\frac{4}{3}\,\frac{\mathbb{E}_{\nu}}{m_{Z}^{2}}\right)\,,$ (58)
$\tilde{H}_{\bar{\nu}}$ $=U\frac{\mathbb{M}^{2}}{2y}U^{\dagger}-\frac{\sqrt{2}G_{\rm F}m_{e}^{4}}{x^{4}}\,\mathbb{N}_{\nu}-\frac{2\sqrt{2}G_{\rm F}ym_{e}^{6}}{x^{6}}\left(\frac{\mathbb{E}_{\ell}+\mathbb{P}_{\ell}}{m_{W}^{2}}+\frac{4}{3}\,\frac{\mathbb{E}_{\nu}}{m_{Z}^{2}}\right)\,.$
Additionally, $\tilde{\mathcal{I}}$ and $\tilde{\bar{\mathcal{I}}}$ represent the comoving collision terms for neutrinos and antineutrinos, respectively. The computation of these collision terms is the most complex and time-consuming aspect of solving the QKEs. Fortunately, the nine-dimensional collision integrals can be simplified to two-dimensional integrals by exploiting the homogeneous and isotropic conditions and the specific forms of the scattering amplitudes Hannestad:1995rs ; Dolgov:1997mb ; Grohs:2015tfy ; Blaschke:2016xxt . The explicit forms of the reduced collision terms are provided in Appendix A.
Furthermore, as noted in the discussion following the QKEs (Eq. (53)), the chemical potential of electrons and the $\mathbb{N}_{l}$ term in the Hamiltonian are negligible due to ue to the very low baryon-to-photon ratio $\eta$ and the efficient electromagnetic interactions. Consequently, the Hamiltonian is predominantly influenced by the net number densities of neutrinos when primordial lepton asymmetries are sufficiently large and vary among different flavors. This results in synchronized flavor oscillations of neutrinos and antineutrinos Bell:1998ds ; Pastor:2001iu ; Abazajian:2002qx ; Wong:2002fa . Consequently, neutrino asymmetries tend to equilibrate among different flavors prior to the epoch of BBN Froustey:2021azz ; Froustey:2024mgf . Since this work focuses on the implications of neutrino asymmetries for BBN, CMB, and LSS, we will concentrate on flavor-equilibrated primordial asymmetries, i.e., the initial primordial asymmetries are set to $\xi_{\nu_{e}}=\xi_{\nu_{\mu}}=\xi_{\nu_{\tau}}=\xi_{n}$, where $\xi_{\nu_{\alpha}}\equiv\mu_{\nu_{\alpha}}/T_{\rm CM}$ is the degeneracy parameter for neutrinos and antineutrinos of flavor $\alpha$. By adopting this initial condition, we bypass the intricate mapping between arbitrary primordial neutrino asymmetries and those after flavor equilibration, leaving the investigation of this mapping for future work. Recent study on such mapping and the implications of distinct neutrino asymmetries among flavors for BBN can be found in  Froustey:2024mgf .
On the other hand, it is well-known that the annihilation of electrons and positrons shortly after neutrino decoupling partially heats neutrinos, leading to a slightly increased $N_{\rm eff}$. To accurately determine $N_{\rm eff}$, finite-temperature QED (FTQED) corrections must be considered Fornengo:1997wa ; Mangano:2001iu ; Bennett:2019ewm . In this work, we address three aspects of FTQED corrections: (1) Corrections to the total energy pressure densities, which primarily affect the Hubble rate; (2) Corrections to the electron mass, which primarily influence the collision terms in the QKEs; (3) Corrections to the continuity equation of the Universe, typically expressed as a differential equation for the dimensionless photon temperature $z$,
$\frac{\mathrm{d}z}{\mathrm{d}x}=\frac{rJ_{2}(r)+G_{1}(r)-\frac{1}{4\pi^{2}z^{3}}\int_{0}^{\infty}\mathrm{d}y\,y^{3}\sum_{\alpha}\left(\frac{\mathrm{d}\varrho_{\alpha\alpha}}{\mathrm{d}x}+\frac{\mathrm{d}\bar{\varrho}_{\alpha\alpha}}{\mathrm{d}x}\right)}{\left[r^{2}J_{2}(r)+J_{4}(r)\right]+G_{2}(r)+\frac{2\pi^{2}}{15}}\,,$ (59)
where $r=x/z$, $\frac{\mathrm{d}\varrho_{\alpha\alpha}}{\mathrm{d}x}$, and $\frac{\mathrm{d}\bar{\varrho}_{\alpha\alpha}}{\mathrm{d}x}$ are the diagonal components of the QKEs for neutrinos and antineutrinos. The explicit forms of the remaining $J_{i}(r)$ and $G_{i}(r)$ functions appearing in the equation, as well as other FTQED corrections, can be found in Appendix B. In practice, this equation should be solved in conjunction with the QKEs, and we consider corrections up to order $\mathcal{O}(e^{3})$ while neglecting logarithmic terms in the FTQED corrections as their effects are minimal Bennett:2019ewm ; Bennett:2020zkv ; further details are provided in Appendix B. Other FTQED corrections, such as those to the interaction vertices, are omitted in this work.
Furthermore, we restrict our analysis to the normal mass hierarchy of neutrinos and neglect any possible CP-violating phase in the transformation matrix between the neutrino flavor and mass bases. Based on the most recent values from the Particle Data Group pdg , we adopt the following neutrino mass differences and mixing angles: $\Delta m_{21}^{2}=7.53\times 10^{-5}\,\text{eV}^{2}$, $\Delta m_{31}^{2}=2.53\times 10^{-3}\,\text{eV}^{2}$, $\sin^{2}\theta_{12}=0.307$, $\sin^{2}\theta_{23}=0.545$, and $\sin^{2}\theta_{13}=0.0218$. The results for neutrino decoupling and BBN are independent of the minimal neutrino mass, provided $m_{1}\ll 0.1\,\text{MeV}$. On the other hand, the resulting CMB anisotropies and LSS of the Universe are sensitive to the choice of $m_{1}$, which will be discussed in more detail in Sections 5 and 6.
3.2 Numerical methods
To solve the QKEs for neutrinos and antineutrinos numerically, we have extended the public code FortEPiaNO Gariazzo:2019gyi ; Bennett:2020zkv , initially designed for solving the QKEs in the Standard Model case, to accommodate primordial neutrino asymmetries. This subsection outlines the key numerical methods employed in the code.
For the momentum grid, the comoving momentum $y$ is discretized on a linear grid with $y_{\rm min}=0.01$, $y_{\rm max}=20$, and a total of $N_{y}=40$ points. The integrals are computed using the Newton-Cotes (NC) formula of order 1. The initial conditions are specified with an initial reduced scale factor $x=0.005$, corresponding to a comoving temperature $T_{\rm CM,in}\simeq 102.2\text{ MeV}$. The initial common temperature for all species is computed by conserving total entropy from an even higher temperature, yielding $z_{\rm in}=2.81\times 10^{-6}$. The initial density matrices are given by
$\varrho(x_{\rm in},y)={\rm diag}\{f_{\nu}^{({\rm in})}(y),f_{\nu}^{({\rm in})}(y),f_{\nu}^{({\rm in})}(y)\},\quad\bar{\varrho}(x_{\rm in},y)={\rm diag}\{f_{\bar{\nu}}^{({\rm in})}(y),f_{\bar{\nu}}^{({\rm in})}(y),f_{\bar{\nu}}^{({\rm in})}(y)\}.$ (60)
where
$f_{\nu}^{({\rm in})}(y)\equiv\frac{1}{e^{(y-\xi_{\nu})/z_{\rm in}}+1}\quad\text{and}\quad f_{\bar{\nu}}^{({\rm in})}(y)\equiv\frac{1}{e^{(y+\xi_{\nu})/z_{\rm in}}+1}\,.$ (61)
Finally, the QKEs are solved together with the energy conservation equation for $z$ using the DLSODA method from the ODEPACK hindmarsh1982odepack Fortran library. The absolute and relative error tolerances are set to $10^{-7}$, ensuring that the numerical errors of the results remain below $10^{-5}$.
3.3 Numerical results of neutrino decoupling
We present the numerical results of neutrino decoupling in this subsection. First, we consider the final energy density of neutrinos, which is typically parameterized by the effective number of relativistic species $N_{\rm eff}$, defined as
$N_{\rm eff}\equiv\frac{87}{8}\left(\frac{11}{4}\right)^{4/3}\frac{\rho_{\nu}}{\rho_{\gamma}}\,,\quad\rho_{\nu}=\sum_{\alpha}\int\frac{d^{3}p}{(2\pi)^{3}}\,p\,(\varrho_{\nu_{\alpha}}+\bar{\varrho}_{\nu_{\alpha}})$ (62)
where $\rho_{\nu}$ represents the total neutrino energy density obtained by solving the QKEs and $\rho_{\gamma}$ is the photon energy density. For the SM in the instantaneous decoupling limit, $N_{\rm eff}$ equals the number of neutrino flavors. However, since the temperature of neutrino decoupling is close to that of electron-positron annihilation, neutrinos are also heated through weak interactions with electrons and positrons, resulting in a slightly larger $N_{\rm eff}$. The most accurate calculation for the SM $N_{\rm eff}$ to date, which includes neutrino oscillations, matter effects, FTQED corrections up to $\mathcal{O}(3)$ order, and neutrino-electron and neutrino-neutrino collision terms, yields $N_{\rm eff}^{\rm SM}=3.0440\pm 0.0002$ Froustey:2020mcq ; Bennett:2020zkv .
Figure 2: Left panel: Results of the neutrino non-instantaneous decoupling correction $\delta N_{\rm eff}^{\rm dec}=N_{\rm eff}-N_{\rm eff}^{\rm inst}$ as a function of the neutrino degeneracy parameter $\xi_{\nu}$. The solid and dashed curves represent the results of the numerical solution and the fitting function Eq. (63), respectively. Right panel: Relative differences in the number densities of electron neutrinos and antineutrinos compared to the SM predictions in the instantaneous decoupling limit, $n_{\nu_{e},\bar{\nu}_{e}}\simeq 56\,\text{cm}^{-3}$.
In the presence of non-zero primordial neutrino asymmetries, $N_{\rm eff}$ receives additional contributions from the neutrino chemical potentials. In the instantaneous decoupling limit, the comoving distribution of neutrinos and antineutrinos of flavor $\alpha$ can be approximated as $f_{\nu_{\alpha},\bar{\nu}_{\alpha}}=\frac{1}{e^{y\,\pm\,\xi_{\nu_{\alpha}}}+1}$. Therefore, the total $N_{\rm eff}$ can be estimated by
$N_{\rm eff}^{\rm inst}\simeq 3+\sum\limits_{\alpha}\left(\frac{30}{7\pi^{2}}\xi_{\nu_{\alpha}}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu_{\alpha}}^{4}\right)=3+3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right),$ (63)
where we have used the condition $\xi_{\nu_{e}}=\xi_{\nu_{\mu}}=\xi_{\nu_{\tau}}=\xi_{\nu}$ in the second equality.
[…]
$N_{\rm eff}(\xi_{\nu})=N_{\rm eff}^{\rm SM}+3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right)+0.0102\,\xi_{\nu}^{2},$ (64)
where $N_{\rm eff}^{\rm dec,SM}=3.0440\pm 0.0002$.
[…]
Figure 3: Density matrices of neutrinos and antineutrinos in the flavor basis (left panel) and mass basis (right panel), normalized to the Fermi-Dirac distribution with $\xi_{\nu}=0$. The gray curves represent the density matrices for neutrinos and antineutrinos with $\xi_{\nu}=0$, while the red and green curves represent those for neutrinos and antineutrinos with non-zero $\xi_{\nu}$, respectively. For each case, the different flavor eigenstates ($\nu_{e}$, $\nu_{\mu}$, and $\nu_{\tau}$) and mass eigenstates ($\nu_{1}$, $\nu_{2}$, and $\nu_{3}$) are indicated by solid, dashed, and dash-dotted curves, respectively. Note that the density matrices for $\nu_{\mu}$ and $\nu_{\tau}$ nearly overlap. The Fermi-Dirac distributions with corresponding $\xi_{\nu}$ are also indicated by the gray dotted curves.
Next we present the resulting density matrices for neutrinos and antineutrinos after decoupling. The density matrices as a function of the comoving momentum $y\equiv p/T_{\rm CM}$ in the flavor basis and the mass basis for several choices of $\xi_{\nu}$ are shown in the left and right panels of Fig. 3 respectively, normalized to the Fermi-Dirac distribution with $\xi_{\nu}=0$. The density matrices of neutrinos and antineutrinos with $\xi_{\nu}=0$ coincide with each other and are represented by the gray curves. Our results also reproduce the previous density matrices obtained in the literature. For each choice of In the case of non-zero $\xi_{\nu}$, the density matrices of neutrinos and antineutrinos no longer coincide and are represented by red and green curves, respectively. For each case we use the solid, dashed and dash-dotted curve to indicate the three flavor eigenstates ($\nu_{e}$, $\nu_{\mu}$ and $\nu_{\tau}$) and the three mass eigenstates ($\nu_{1}$, $\nu_{2}$ and $\nu_{3}$), respectively. For each choice of the $\xi_{\nu}$, the corresponding thermal equilibrium distributions for neutrinos and antineutrinos $f_{\rm FD}^{\nu,\bar{\nu}}=1/(e^{y\,\pm\,\xi_{\nu}}+1)$ are also plotted by the dotted curves. Clearly, the density distributions of neutrinos and antineutrinos for the given $\xi_{\nu}$ coincide with the corresponding Fermi-Dirac distributions when $y$ is small, while the spectra are distorted at large $y$. This is due to the fact that the weak interaction rate between neutrinos and electrons/positrons is proportional to the neutrino momentum, so that the neutrino with smaller momentum decouples earlier, preserving the thermal Fermi-Dirac distribution, while the neutrino with larger momentum decouples later, resulting in a larger spectral distortion with respect to the Fermi-Dirac distribution. Regarding the density distributions for different flavors, we see that electron neutrinos possess larger spectral distortions due to the additional charge-current weak interaction, while the spectral distortions of muon neutrinos almost coincide with those of tauon neutrinos, since they have the same weak interactions. Instead, the spectral distortions of the three neutrino mass eigenstates are uniformly separated, as shown in the right panel.
Figure 4: The spectral distortions of neutrinos and antineutrinos in the flavor basis (top left) and mass basis (lower left) with respect to the Fermi-Dirac distribution with corresponding $\xi_{\nu}$. The meaning of the curves is the same as in the previous plot Fig. 3, while the spectral distortions of tauon neutrino in the left panel are not shown for brevity. The zoomed version of the spectral distortions in the flavor and mass basis are shown in the right panel.
Considering that the density distributions of neutrinos and antineutrinos coincide with the corresponding thermal Fermi-Dirac distributions for small momentum, it is natural to define the spectral distortions as $\delta\varrho_{\nu,\bar{\nu}}\equiv\delta\varrho_{\nu,\bar{\nu}}/f_{FD}(\pm\xi)-1$, where the plus sign corresponds to neutrinos and the minus sign to antineutrinos. The resulting spectral distortions as a function of the momentum $y$ in the flavor and mass basis are presented in the upper left and upper right panels of Fig. 4, respectively. The meaning of the curves is the same as in the previous plot Fig. 3, while the spectral distortions of the tauon neutrinos are not shown in the left panels, since they almost coincide with the spectral distortions of the muon neutrinos. We also plot the zoomed version of the spectral distortions in the lower panels. Interestingly, the spectral distortions for neutrinos and antineutrinos with non-zero $\xi_{\nu}$ differ from each other and lie on different sides of the spectral distortions with $\xi_{\nu}=0$. The antineutrinos with non-zero $\xi_{\nu}$ possess a relatively larger spectral distortion than that with $\xi_{\nu}=0$, while the neutrinos possess a relatively smaller spectral distortion. In addition, the larger the $\xi_{\nu}$, the larger the discrepancy between the spectral distortions of neutrinos and antineutrinos, clearly seen from the zoomed plots. Such properties are both valid in the flavor basis and in the mass basis. Finally, the above conclusions about neutrino density matrices also hold for negative $\xi_{\nu}$, except that the density matrices for neutrinos and antineutrinos with $-\xi_{\nu}$ are exchanged comparing with those for $\xi_{\nu}$. In other words, the density distributions and spectral distortions for antineutrinos in Fig. 3 and Fig. 4 are also the density distributions and spectral distortions for neutrinos with the opposite $\xi_{\nu}$.
As a brief conclusion of this section, we have presented the most accurate results of neutrino decoupling in the presence of primordial neutrino asymmetries and discussed the resulting $N_{\rm eff}$ and the spectral distortions of neutrinos and antineutrinos. In the next sections, we will study the effects of primordial neutrino asymmetries on the processes of BBN, CMB and structure formation, while emphasizing the corrections induced by the spectral distortions of neutrinos and antineutrinos.
4 Impacts of the Primordial Neutrino Asymmetry on BBN
Soon after neutrino decoupling, there is an epoch of primordial nucleosynthesis or the so-called Big Bang Nucleosynthesis (BBN), where the light elements are produced after protons and neutrons decouple from the SM thermal bath. Since the decoupling of protons depends on electron neutrinos, the resulting abundances of light elements after BBN depends on the properties of electron neutrinos and antineutrinos. BBN is affected by primordial neutrino asymmetries in three ways as follow.
  1. 1.
    The leading effect is the change in the weak interaction rate at the onset of BBN due to the chemical potential of the electron neutrinos, since the conversion between proton and neutron is related to the electron neutrino through the weak interaction, so that the difference of the chemical potentials of proton and neutron is decided by the chemical potential of electron neutrino, $i.e.,$ $\mu_{n}-\mu_{p}=\mu_{e}-\mu_{\nu_{e}}\simeq-\mu_{\nu_{e}}$. Therefore, the resulting neutron-to-proton ratio and the subsequent BBN processes are modified according to the absolute value of the electron neutrino chemical potential and its sign.
  2. 2.
    In addition, the expansion rate of the universe during BBN is changed by the additional neutrino energy density contributed by the chemical potentials. As we have seen in the last section, the neutrino energy density after neutrino decoupling is usually represented by $N_{\rm eff}$, and we have $N_{\rm eff}\simeq 3.044+3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right)$. The corresponding correction in BBN therefore depends on the absolute asymmetries of all three neutrino flavors.
  3. 3.
    Beyond the approximation of the thermal Fermi-Dirac distributions for neutrinos and antineutrinos, the spectral distortions resulting from the non-instantaneous decoupling of the neutrinos will also induce a correction of the weak interaction rate and thus affect the BBN. Such corrections are usually neglected in previous studies, especially in the presence of primordial neutrino asymmetries.
In the following, we first review the general treatment of the weak interaction rate and introduce the new correction induced by the neutrino spectral distortions, then we discuss the resulting light element abundances from the BBN.
4.1 The weak correction from neutrino spectral distortions
At the onset of the BBN, the conversion between neutron ($n$) and proton ($p$) is dominated by following interaction reactions,
$n+\nu_{e}$ $\leftrightarrow p+e^{-},$ (65)
$n$ $\leftrightarrow p+e^{-}+\bar{\nu}_{e},$ (66)
$n+e^{+}$ $\leftrightarrow p+\bar{\nu}_{e}.$ (67)
In the remainder of this section, we will omit the subscript for the electron-neutrino and the superscript for the electron/positron as long as there is no ambiguity. Due to the forceful coupling between photons and electrons and positrons, it is appropriate to assume that all particles expect for neutrinos are in local thermal equilibrium. Therefore, the neutron and proton number densities obey the Boltzmann equations
$\dot{n}_{n}+3Hn_{n}$ $=-n_{n}\Gamma_{n\to p}+n_{p}\Gamma_{p\to n},$ (68)
$\dot{n}_{p}+3Hn_{p}$ $=-n_{p}\Gamma_{p\to n}+n_{n}\Gamma_{n\to p},$ (69)
where the reaction rates $n_{n}\Gamma_{n\to p}$ and $n_{p}\Gamma_{p\to n}$ incorporate all relevant collision terms
$\Gamma_{n\to p}$ $=$ $\Gamma_{n+\nu\to p+e}+\Gamma_{n\to p+e+\bar{\nu}}+\Gamma_{n+e\to p+\bar{\nu}},$ (70)
$\Gamma_{p\to n}$ $=$ $\Gamma_{p+e\to n+\nu}+\Gamma_{p+e+\bar{\nu}\to n}+\Gamma_{p+\bar{\nu}\to n+e}.$ (71)
And the reaction rate for a specific interaction reaction with the nucleon $a$ in the initial state and nucleon $b$ in the final state takes the form as
$n_{a}\Gamma_{a\to b}=$ $\int\mathrm{d}\pi_{a}\mathrm{d}\pi_{b}\mathrm{d}\pi_{e}\mathrm{d}\pi_{\nu}\delta^{(4)}\left({p}_{a}-p_{b}+\alpha_{e}p_{e}+\alpha_{\nu}p_{\nu}\right)\left|M\right|^{2}_{a\to b}{}$ (72)
×$\times f_{a}(E_{a})f_{b}(-E_{b})f_{e}(\alpha_{e}{E}_{e})\tilde{f}_{\nu}(E_{\nu}),$
where $\mathrm{d}\pi_{i}=\frac{1}{(2\pi)^{3}}\frac{\mathrm{d}^{3}\mathbf{p}_{i}}{2E_{i}}$, $\alpha_{e}(\alpha_{\nu})=1$ if the electron (neutrino) is in the initial state and $\alpha_{e}(\alpha_{\nu})=-1$ if the electron (neutrino) is in the final state. In addition, $f_{i}(E)\equiv\frac{1}{e^{E/T_{\gamma}}+1}$ for $i=a,b,e$ is the Fermi-Dirac distribution with vanishing chemical potential, and we have used the relation $1-f_{i}(E)=f_{i}(-E)$ to incorporate the corresponding Pauli blocking factors. Lastly, $\tilde{f}_{\nu}(E_{\nu})$ equals $\varrho_{\nu}(E_{\nu})\,(\varrho_{\bar{\nu}}(E_{\nu}))$ for the neutrino (antineutrino) in the initial state and equals $1-\varrho_{\nu}(E_{\nu})\,(1-\varrho_{\bar{\nu}}(E_{\nu}))$ for the neutrino (antineutrino) in the initial state.
[…]
$\bar{\Gamma}_{n\to p}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p[\chi_{+}(E)+\chi_{+}(-E)]\,,$ (73)
$\bar{\Gamma}_{p\to n}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p[\chi_{-}(E)+\chi_{-}(-E)]\,,$ (74)
with $E=\sqrt{p^{2}+m_{e}^{2}}$ is the energy of the electron and $K=1/(\tau_{n}\lambda_{0}m_{e}^{5})$, where $\tau_{n}$ is the life time of the neutron and $\lambda_{0}$ the Born coefficient. The $\chi^{\pm}$ function is defined as
$\chi_{\pm}(E)\equiv(E_{\nu}^{\mp})^{2}f_{e}(-E)f_{\nu}^{\pm}(E^{\mp}_{\nu})\,,\,\,E^{\mp}_{\nu}\equiv E\mp\Delta,$ (75)
where $\Delta\equiv m_{n}-m_{p}$ is the energy gap between neutron and proton. The first term in Eq. (73) corresponds to the $n\to p+e$ processes Eq. (65) or Eq. (66) depending on whether $E-\Delta$ is positive or negative, and the second term corresponds to the $n+e\to p$ process Eq. (67) and similarly for the terms in Eq. (74). Note that here $E^{\mp}_{\nu}$ is equal to the energy of the neutrino $E_{\nu}$ for the processes where the neutrino is in the initial state, or its opposite $-E_{\nu}$ for the processes where the neutrino is in the final state. And the term $f_{\nu}^{\pm}(E^{\mp}_{\nu})$ in the function $\chi^{\pm}$ gives the corresponding distribution or the Pauli-blocking factor by the relation $1-f_{\nu}^{+}(E_{\nu})=f_{\nu}^{-}(-E_{\nu})$.
Now we introduce the corrections induced by the spectral distortions of neutrinos and antineutrinos. In the infinite nucleon mass limit, the corrections to the Born rates can be written as
$\delta\bar{\Gamma}_{n\to p}^{\rm SD}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p[\delta\chi_{+}(E)+\delta\chi_{+}(-E)]\,,$ (76)
$\delta\bar{\Gamma}_{p\to n}^{\rm SD}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p[\delta\chi_{-}(E)+\delta\chi_{-}(-E)]\,,$ (77)
with
$\delta\chi_{\pm}(E)\equiv$ $(E_{\nu}^{\mp})^{2}f_{e}(-E){\rm sgn}(E^{\mp}_{\nu}){}$ (78)
×$\times\left(\frac{1+\Theta(\pm E^{\mp}_{\nu})\delta\varrho_{\nu}+\Theta(\mp E^{\mp}_{\nu})\delta\varrho_{\bar{\nu}}}{e^{\left|E^{\mp}_{\nu}\right|/T_{\rm CM}\mp{\rm sgn}(E^{\mp}_{\nu})\xi_{\nu}}+1}-\frac{1}{e^{\left|E^{\mp}_{\nu}\right|/T_{\nu}\mp{\rm sgn}(E^{\mp}_{\nu})\xi_{\nu}}+1}\right)\,,$
where $\Theta(x)$ is the Heaviside step function, $\delta\varrho_{\nu}$ and $\delta\varrho_{\bar{\nu}}$ are the spectral distortions of electron neutrinos and antineutrinos. For a particular process, the various sgn functions and the $\Theta$ functions together represent the nature and the position of the neutrino particle: Whether it is a neutrino or an antineutrino, and whether it is in the initial state or the final state. We also note that $T_{\rm CM}$ and $T_{\nu}$ has slight difference after neutrino decoupling. If we neglect such difference and set $\xi_{\nu}=0$, so that $\delta\varrho_{\nu}=\delta\varrho_{\bar{\nu}}$, then Eq. (78) coincides with Eq. (29) in Ref. Froustey:2019owm , where they considered the corrections to the weak rates from neutrino spectral distortions in the SM.
Finally, we briefly review the other corrections to the weak interaction rates that we have considered in this work. First are the zero temperature radiative corrections, including the so-called pure radiative corrections corresponding to the emission and absorption of a virtual photon in the interaction and the so-called bremsstrahlung correction due to the emission of a real photon. Second are the finite temperature radiative corrections due to the finite temperature corrections to the proton-electron interactions and the electron energy. Third and fourth are the corrections from the finite nucleon mass and the weak magnetism. Combining all these corrections, the total weak reaction rates are given by
$\Gamma_{n\to p}=\Gamma^{\rm RC0}_{n\to p}+\delta\Gamma^{\rm SD,RC0}_{n\to p}+\Gamma^{T}_{n\to p}+\delta\Gamma^{\rm BS}_{n\to p}+\delta\Gamma^{\rm FM}_{n\to p}\,,$ (79)
where the superscript ’RC0’, ’SD’, ’T’, ’BS’ and ’FM’ represent the zero temperature radiative correction, the spectral distortion, the finite temperature radiative correction, the bremsstrahlung and the finite nucleon mass correction, respectively. And the form for the $p\to n$ rates is similar. Note that the effects of different corrections may be coupled. For example, the effects of the zero temperature radiative correction on the corrections induced by the spectral distortions are also considered in this work. Specifically, the Born rates with the zero temperature radiative correction take the form of
${\Gamma}^{{\rm RC}0}_{n\to p}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p[\mathcal{F}_{+}(E)R(E,|\Delta-E|)\chi_{+}(E)+\mathcal{F}_{+}(-E)R(E,\Delta+E)\chi_{+}(-E)]\,,$ (80)
${\Gamma}^{{\rm RC}0}_{p\to n}$ $=K\int_{0}^{\infty}p^{2}\mathrm{d}p[\mathcal{F}_{-}(E)R(E,\Delta+E)\chi_{-}(E)+\mathcal{F}_{-}(-E)R(E,|E-\Delta|)\chi_{-}(-E)]\,.$ (81)
The explicit expressions for the Fermi function $F_{\pm}$ and the radiative correction factor $R(E,k_{\rm max})$ can be found in $e.g.,$   Ref. Pitrou:2018cgg and are not repeated here for brevity. We see that the zero temperature radiative correction only modifies the coefficients of the $\chi_{\pm}$ function. The $\delta\Gamma^{\rm RC0,SD}_{n\to p}$ and $\delta\Gamma^{\rm RC0,SD}_{p\to n}$ can also be obtained simply by the replacement of $\chi_{\pm}\to\delta\chi_{\pm}$. In principle, the other corrections are also affected by the spectral distortions of neutrinos and antineutrinos. However, such effects must be negligible since the corrections themselves are already sub-leading. Therefore, we neglect these interferences in this work.
4.2 Results of BBN abundances
In this subsection we present the resulting light element abundances incorporating all the corrections discussed in the last subsection. To do this, we have added the aforementioned correction induced by the neutrino spectral distortions in the public BBN code PRIMATPitrou:2018cgg and solved the abundances of the light elements. For the computation we have used the latest values of other physical and cosmological parameters, $e.g.,$ the neutron lifetime is taken as $\tau_{n}=879.4s$, the axial coupling of the nucleons is $g_{A}=1.2756$, and the baryon density is $\omega_{b}h^{2}=0.024$. For convenience, we define the mass fraction of isotope $i$ as $X_{i}\equiv A_{i}\frac{n_{i}}{n_{b}}$, where $A_{i}$ and $n_{i}$ are the baryon number and the number density of isotope $i$ and $n_{b}$ is the total number density of baryon. Customarily, we also denote $Y_{P}\equiv X_{{}^{4}{\rm He}}$ and $i/\rm H\equiv n_{i}/n_{{}^{1}\rm H}$.
Figure 5: Light element abundances as a function of the neutrino degeneracy parameter $\xi_{\nu}$, normalized to the results for $\xi_{\nu}=0$. The ${}^{3}{\rm He}$ stands for ${}^{3}{\rm He}+\rm T$ and ${}^{7}{\rm Li}$ stands for ${}^{7}{\rm Li}+\rm^{7}Be$ due to the slow radioactive decays. All the optional weak rate corrections in PRIMAT (see Ref. Pitrou:2018cgg ) and the new correction from neutrino spectral distortions are included in the calculation. The $68\%$ constraints on the helium abundance $Y\rm_{P}$ from the CMB (${Y\rm_{P}}=0.242\pm 0.012$) and the recent EMPRESS experiment (${Y\rm_{P}}=0.2370^{+0.0034}_{-0.0033}$) are indicated by the light and dark blue shaded regions, respectively.
We show in Fig. 5 the resulting light element abundances, including all the aforementioned weak rate corrections in (79) and the QED corrections up to $\mathcal{O}(e^{3})$ for plasma thermodynamics and certain nuclear rates, as a function of the neutrino degeneracy parameter $\xi_{\nu}$. The abundance of ${}^{3}{\rm He}$ stands for ${}^{3}{\rm He}+\rm T$ and ${}^{7}{\rm Li}$ for ${}^{7}{\rm Li}+\rm^{7}Be$ to account for the slow radioactive decays. Since the abundances for different elements can vary by several orders of magnitude, we normalize the element abundances by their corresponding abundance with $\xi_{\nu}=0$. Therefore all curves in Fig. 5 intersect at $\xi_{\nu}=0$. We see that the abundances of both $\rm D,\,^{3}{\rm He},\,^{4}{\rm He}$ and ${}^{7}{\rm Li}$ decrease with increasing $\xi_{\nu}$ due to the fact that the larger the $\xi_{\nu}$, the smaller the number density ratio between neutron and proton. In addition, the variation of the ${}^{4}{\rm He}$ abundance is the most sensitive to the variation of $\xi_{\nu}$, while the ${}^{3}{\rm He}$ abundance is the least sensitive. We also show in Fig. 5 the constraints on the helium abundance from obtained by CMB and EMPRESS experiments,
${Y\rm_{P}}|_{\rm CMB}=0.242\pm 0.012,\,\,{Y\rm_{P}}|_{\rm EMPRESS}=0.2370^{+0.0034}_{-0.0033},$ (82)
by the light and dark blue shaded regions. From our results, the constraints on ${Y\rm_{P}}$ within the $68\%$ confidence interval can be easily translated into constraints on the neutrino degeneracy parameter $\xi_{\nu}$ as
${\rm CMB}:\,\,-0.026\leq$ $\,\xi_{\nu}\leq 0.069\,,$ (83)
EMPRESS:  ${\rm EMPRESS}:\,\,0.027\leq$ $\,\xi_{\nu}\leq 0.054\,.$ (84)
Clearly, the EMPRESS experiment prefer a nonzero positive $\xi_{\nu}$ while the CMB experiment is still compatible with the SM case $\xi_{\nu}=0$.
Figure 6: The relative difference of the light element abundances compared to the results in the neutrino instantaneous decoupling limit as a function of the neutrino degeneracy parameter $\xi_{\nu}$. The solid and dotted curves correspond to the results with and without the weak corrections induced by the neutrino spectral distortions, respectively. The meaning of the color for the curves is the same as in the plot Fig. 5, except that the negative results of ${}^{7}{\rm Li}$ are shown here for simplicity.
BBN framework ${Y\rm_{P}}$ ${\rm D}/{\rm H}\times 10^{5}$ ${{}^{3}{\rm He}}/{\rm H}\times 10^{5}$ ${{}^{7}{\rm Li}}/{\rm H}\times 10^{10}$
Inst.dec, $\xi_{\nu}=-0.1$ 0.27359 2.58214 1.06094 5.87974
Inst.dec, $\xi_{\nu}=0$ 0.24711 2.42833 1.03799 5.53146
Inst.dec, $\xi_{\nu}=0.1$ 0.22260 2.30372 1.01897 5.14551
No-SD, $\xi_{\nu}=-0.1$ 0.27375 2.59246 1.06239 5.85555
No-SD, $\xi_{\nu}=0$ 0.24712 2.43725 1.03926 5.5077
No-SD, $\xi_{\nu}=0.1$ 0.2225 2.31162 1.02011 5.12186
Actual-SD, $\xi_{\nu}=-0.1$ 0.27375 2.59247 1.06239 5.85558
Actual-SD, $\xi_{\nu}=0$ 0.24721 2.43769 1.03933 5.50905
Actual-SD, $\xi_{\nu}=0.1$ 0.22266 2.31244 1.02024 5.12449
Table 1: Light element abundances including all weak rate corrections in (79), for various implementations of neutrino decoupling and degeneracy parameters. ${}^{3}{\rm He}$ stands for $(^{3}{\rm He}+\mathrm{T})$ and ${}^{7}{\rm Li}$ stands for $(^{7}{\rm Li}+\rm Be)$ to account for slow radioactive decays.
BBN framework $\delta{Y\rm_{P}}\,(\%)$ $\delta{\rm D}/{\rm H}\,(\%)$ $\delta{{}^{3}{\rm He}}/{\rm H}\,(\%)$ $\delta{{}^{7}{\rm Li}}/{\rm H}\,(\%)$
No-SD, $\xi_{\nu}=-0.1$ 0.058 0.4 0.137 -0.411
No-SD, $\xi_{\nu}=0$ 0.007 0.367 0.122 -0.43
No-SD, $\xi_{\nu}=0.1$0.048 0.343 0.112 -0.46
Actual-SD, $\xi_{\nu}=-0.1$ 0.059 0.4 0.137 -0.411
Actual-SD, $\xi_{\nu}=0$ 0.042 0.385 0.129 -0.405
Actual-SD, $\xi_{\nu}=0.1$ 0.026 0.379 0.125 -0.408
Table 2: Relative difference of the light element abundances for the case with and without the corrections induced by the neutrino spectral distortions compared to the results in the neutrino instantaneous decoupling limit. ${}^{3}{\rm He}$ stands for $(^{3}{\rm He}+\mathrm{T})$ and ${}^{7}{\rm Li}$ stands for $(^{7}{\rm Li}+\rm Be)$ to account for slow radioactive decays.
It is also interesting to discuss the effects of the corrections induced by the neutrino spectral distortions on BBN. Focusing on the corrections induced by different treatments of the neutrino decoupling process, we consider the following three implementations: (1) The neutrino instantaneous decoupling limit, $i.e.,$ the $N_{\rm eff}^{\rm inst}$ as discussed in the last section and the thermal Fermi-Dirac distributions for neutrinos and antineutrinos are used in the calculation of the BBN; (2) The no-SD case, where the actual $N_{\rm eff}$ and neutrino temperatures are used, including the non-instantaneous decoupling effect, but considering only the thermal Fermi-Dirac distributions for neutrinos and antineutrinos. This is the most common treatment for BBN in the studies to date; (3) The actual SD case, using the actual $N_{\rm eff}$ with the non-instantaneous decoupling effect and the actual density distributions with the spectral distortions for neutrinos and antineutrinos, corresponding to the results shown in Fig. 5. Using the BBN results in the instantaneous decoupling limit as the benchmark, we show the relative differences of the results for the no-SD and the actual-SD cases with the dotted and solid curves in Fig. 6, respectively. Note that we have plotted the opposite of the relative differences for ${}^{7}{\rm Li}$ in Fig. 6 as it is negative due to the particular nuclear reaction rates used in PRIMAT. We also report the explicit values of the abundances of helium-4, deuterium, helium-3 and lithium-7 for the degeneracy parameters $\xi_{\nu}=0,\pm 1$ in table 1, and the associated relative variations compared to the baseline abundance in table 2. We also note that our results with $\xi_{\nu}=0$ are consistent with the results reported in Froustey:2020mcq , where they also considered the effects of neutrino spectral distortions to BBN within the SM case.
As we can see in Fig. 6 and the tables, the actual-SD case and the no-SD case yield almost the same light element abundances for $\xi_{\nu}=-0.1$. And the variation of the BBN abundances from the results in the instantaneous decoupling limit for the actual-SD case show little dependence on $\xi_{\nu}$, while the traditional treatment without the spectral distortions predicts a stronger dependence for both elements. In particular, the no-SD case predicts negative variations for the helium abundance with positive $\xi_{\nu}$ compared to the result in the instantaneous decoupling limit. However, the more precise results with the spectral distortions show that the variations of the helium abundance are positive, showing that the inclusion of the neutrino spectral distortions is necessary for a precise and accurate prediction of the BBN abundances.
Finally, we also provide a set of fitting functions for the light element abundances as
$Y_{\mathrm{P}}\left(\xi_{\nu}\right)$ $\simeq\left.Y_{\mathrm{P}}\right|_{\xi_{\nu}=0}\times e^{-1.033\xi-0.131\xi^{2}},\,\,$ (85)
$\mathrm{D/H}\left(\xi_{\nu}\right)$ $\simeq\left.\mathrm{D/H}\right|_{\xi_{\nu}=0}\times e^{-0.57\,\xi_{\nu}(1-\xi_{\nu})}\,,$ (86)
${}^{3}\mathrm{He/H}\left(\xi_{\nu}\right)$ $\simeq\left.{}^{3}\mathrm{He/H}\right|_{\xi_{\nu}=0}\times e^{-0.19\,\xi_{\nu}(1-\xi_{\nu})},\,\,$ (87)
${}^{7}\mathrm{Li/H}\left(\xi_{\nu}\right)$ $\simeq\left.{}^{7}\mathrm{Li/H}\right|_{\xi_{\nu}=0}\times e^{-0.63\,\xi_{\nu}(1+\xi_{\nu})},$ (88)
where the subscript $\xi_{\nu}=0$ labels the results obtained in the SM case $\xi_{\nu}=0$ using the actual density matrices, see the explicit values in Table 1. Such fitting functions benefit for further studies, and the fractional errors are smaller than $0.1\%$ with the range $-0.1\leq\xi_{\nu}\leq 0.1$.  [^3] The coefficients of our fitting functions are slightly different from the results in Pitrou:2018cgg ; Froustey:2024mgf , which we think is due to the different corrections to the weak interaction rates we considered.
5 Impacts of the Primordial Neutrino Asymmetry on CMB and LSS
In this section we focus on the impacts of primordial neutrino asymmetries on the evolution of the late universe, especially the cosmic microwave background (CMB) and the large scale structure (LSS). In general, the primordial neutrino asymmetry influence CMB and LSS in three ways.
  1. 1.
    The expansion rate of the Universe is changed by the increased $N_{\rm eff}$ due to the primordial neutrino asymmetries. Particularly, the time of matter-radiation equality is delayed by the additional radiation density, leading to an enhanced early integrated Sachs-Wolfe effect for CMB angular power spectrum.
  2. 2.
    As we have seen in Fig. 5, the abundance of helium is significantly modified by the electron neutrino asymmetry, which will in turn alter the evolution of the late Universe, such as the tail of the CMB angular power spectrum by the amount of diffusion damping and the baryon acoustic oscillation (BAO).
  3. 3.
    In addition, in order to incorporate the effect of neutrino oscillation in the process of neutrino decoupling, the mass of neutrinos must be considered. The neutrino masses will alter the evolution of the late universe by such as, affect the matter-radiation equality, even in the SM case with the thermal Fermi-Dirac distributions. For the SM case, the experiment on CMB and BAO can put a relative stringent bound on the sum of neutrino masses $\sum m_{\nu}<0.07\,{\rm eV}$ DESI:2024mwx , we will return to the constraint on neutrino masses in the next section. Here we emphasize that the neutrino Boltzmann hierarchy for massive neutrinos will be furthermore modified by the neutrino degeneracy parameter and the spectral distortions.
In the following we first review the impact of massive neutrinos on the universe and present the Boltzmann hierarchies of electrons, photons and neutrinos. Then we discuss the effects of the primordial neutrino asymmetries on the CMB temperature anisotropies and the linear matter power spectrum computed with the Einstein-Boltzmann code CLASS Lesgourgues:2011re ; Lesgourgues:2011rh .
5.1 Effects of the primordial neutrino asymmetries in structure formation
To describe the flavor oscillations in the neutrino decoupling process, the mass of the neutrinos must be taken into account, which in turn has significant implications for the process of structure formation, since massive neutrinos are the only species in the SM that scale as radiation in the early universe and as matter in the late universe. In the early universe, neutrinos and antineutrinos remain relativistic after decoupling as long as their effective temperatures are much larger than their masses. In this case, the neutrino and antineutrino energy densities contribute to the total radiation energy density, parameterised by the aforementioned $N_{\rm eff}$. Since the neutrino effective temperatures decrease as $a^{-1}$ with the expansion of the universe, neutrinos and antineutrinos become non-relativistic when their effective temperatures fall below their masses, and contribute to matter after that. The non-relativistic transitions and the free-steaming effects of massive neutrinos then leave significant imprints on the CMB anisotropies and the LSS. As mentioned above, the situation is modified in the presence of the primordial neutrino asymmetries. In addition to the indirect change induced by the modified Hubble rate and the helium abundance, we highlight several other modifications due to the primordial neutrino asymmetries in the following.
5.1.1 Effects on neutrino abundance
First, the total relic abundance of neutrinos and antineutrinos are directly increased due to the primordial neutrino asymmetries. At present, the total neutrino relic abundance can be written as
$\Omega_{\nu}h^{2}$ $=\frac{\sum\limits_{i}T_{\rm CM,0}^{4}\int_{0}^{\infty}\mathrm{d}yy^{2}\sqrt{y^{2}+m_{i}}\left(\varrho_{i}(y)+\bar{\varrho}_{i}(y)\right)/(2\pi^{2})}{\rho_{\rm crit}}$ (89)
$\overset{\delta\varrho=0}{\simeq}\frac{\sum\limits_{i}m_{i}}{93.12\,{\rm eV}}\frac{{\rm Li}_{3}(-\exp(-\xi_{\nu}))+{\rm Li}_{3}(-\exp(\xi_{\nu}))}{3\zeta(3)/2},$
where $y=p/T_{\rm CM}$ is the comoving momentum defined in section 3, $T_{\rm CM,0}$ is the current comoving temperature, $\rho_{\rm crit}$ is the critical energy density, $\varrho_{i}(y)$ and $\bar{\varrho}_{i}(y)$ are the respective comoving density distributions for neutrinos and antineutrinos with $i=1,2,3$, corresponding to the three mass eigenstates with mass $m_{i}$. Obviously, the present neutrino abundance is also modified by the spectral distortions of neutrinos and antineutrinos, and the integrals over the neutrino distributions cannot be carried out analytically in this case. The effect of the primordial neutrino asymmetries will be clearer by neglecting the spectral distortions and performing the integral analytically, as seen in the second row of (89). The first factor in the second row $\frac{\sum\limits_{i}m_{i}}{93.12\,{\rm eV}}$ is the usual neutrino relic abundance in the SM case, while the second factor represents the effect of neutrino asymmetries with ${\rm Li}_{j}(x)$ the polylogarithms of order $j$ and $\zeta(x)$ the Riemann zeta function. For the SM case one finds ${\rm Li}_{3}(-1)=3\zeta(3)/4$, so the second factor becomes 1. For non-zero $\xi_{\nu}$, the second factor depends only on the absolute value of $\xi_{\nu}$, and the larger the $\xi_{\nu}$, the larger the second factor and the larger the neutrino abundance. However, we note that such corrections from primordial neutrino asymmetries are quite small. For example, the resulting relative difference of $\Omega_{\nu}h^{2}$ for $\xi_{\nu}=0.2$ is only $1.54\%$. Similarly, the corrections to the neutrino abundance induced by the spectral distortions of neutrinos and antineutrinos are also negligible.
5.1.2 Effects on the neutrino Boltzmann hierarchies
In order to obtain the accurate CMB spectral power spectrum and the matter power spectrum, we need to solve a coupled set of the Boltzmann-Einstein equations for the perturbations of the metric, photon, baryon, cold dark matter and massive neutrinos. Here we briefly review how neutrino asymmetries and spectral distortions affect the Boltzmann-Einstein equations. Following Ref. Ma:1995ey , the perturbed distribution functions for the three mass eigenstate neutrinos can be written as
$\varrho_{i}\left(\vec{x},\vec{p},\tau\right)=\varrho_{i}^{0}(y)\left[1+\Psi_{i}\left(\vec{x},y,\hat{n},\tau\right)\right]\,,$ (90)
[…]
$ds^{2}=a^{2}(\tau)\{-d\tau^{2}+(\delta_{ij}+h_{ij})dx^{i}dx^{j}\}.$ (91)
The perturbation of the metric $h_{ij}$ can be decomposed as $h_{ij}=h\delta_{ij}+h_{ij}^{\parallel}+h_{ij}^{\perp}+h_{ij}^{T}$, where $h\equiv h_{ii}$ is the trace part of $h_{ij}$ , and $h_{ij}^{\parallel}$, $h_{ij}^{\perp}$ and $h_{ij}^{T}$ are the traceless parts, satisfying
$\epsilon_{ijk}\partial_{j}\partial_{l}h_{lk}^{\parallel}=0,\quad\partial_{i}\partial_{j}h_{ij}^{\perp}=0,\quad\partial_{i}h_{ij}^{T}=0.$ (92)
By definition, $h_{ij}^{\parallel}$ and $h_{ij}^{\perp}$ can be written in terms of a scalar filed $\mu$ and a divergenceless vector $\vec{A}$,
$h_{ij}^{\parallel}$ $=\left(\partial_{i}\partial_{j}-\frac{1}{3}\delta_{ij}\nabla^{2}\right)\mu$ (93)
$h_{ij}^{\perp}$ $=\partial_{i}A_{j}+\partial_{j}A_{i},\quad\partial_{i}A_{i}=0.$
Therefore, the scalar modes of the metric perturbations are characterised by the two scalar fields $h$ and $\mu$, while the vector and the tensor modes are characterised by $\vec{A}$ and $h_{ij}^{T}$, respectively.
In this work we will only consider the scalar modes of $h_{ij}$, which can be Fourier transformed as
$h_{ij}(\vec{x},\tau)=\int d^{3}k~{}e^{i\vec{k}\cdot\vec{x}}\left\{\frac{{k}_{i}{k}_{j}}{k^{2}}h(\vec{k},\tau)+6\eta(\vec{k},\tau)\left(\frac{{k}_{i}{k}_{j}}{k^{2}}-\frac{1}{3}\delta_{ij}\right)\right\}\,,$ (94)
where $h(\vec{k},\tau)$ and $\eta(\vec{k},\tau)$ are the trace part and the traceless part of the scalar modes in the Fourier space which are related to the scalar fields $h$ and $\mu$ in the real space. Since neutrinos are collisionless after decoupling, the evolution of $\Psi_{i}$ obeys the first-order collsionless Boltzmann equation in the Fourier-space
$\frac{\partial\Psi_{i}}{\partial\tau}+i\frac{q}{\epsilon}(\vec{k}\cdot\hat{n})\Psi_{i}+\frac{d\ln\varrho_{i}^{0}}{d\ln q}\left[\dot{\eta}-\frac{\dot{h}+6\dot{\eta}}{2}(\hat{k}\cdot\hat{n})^{2}\right]=0\,,$ (95)
where $\epsilon\equiv aE=\sqrt{q^{2}+a^{2}m^{2}_{\nu_{i}}}$ is the comoving energy. Since Eq. (95) is independent of the azimuth angle, we can expand $\Psi$ in a Legendre series
$\Psi_{i}(\vec{k},\hat{n},q,\tau)=\sum_{l=0}^{\infty}(-i)^{l}(2l+1)\Psi_{i,l}(k,q,\tau)P_{l}(\hat{k}\cdot\hat{n})\,.$ (96)
Inserting Eq. (96) into Eq. (95) and using the orthonormality of the Legendre polynomials, we can obtain the Boltzmann hierarchies for $\Psi_{i,l}$:
$\dot{\Psi}_{i,0}$ $=-\frac{qk}{\epsilon}\Psi_{1}+\frac{\dot{h}}{6}\frac{{d}\ln\varrho_{i}^{0}}{{d}\ln q}\,,$ (97)
$\dot{\Psi}_{i,1}$ $=\frac{qk}{3\epsilon}\left(\Psi_{0}-2\Psi_{2}\right)\,,$
$\dot{\Psi}_{i,2}$ $=\frac{qk}{5\epsilon}\left(2\Psi_{1}-3\Psi_{3}\right)-\left(\frac{\dot{h}}{15}+\frac{2\dot{\eta}}{5}\right)\frac{{d}\ln\varrho_{i}^{0}}{{d}\ln q}\,,$
$\dot{\Psi}_{i,l\geq 3}$ $=\frac{qk}{(2l+1)\epsilon}\left[l\Psi_{l-1}-(l+1)\Psi_{l+1}\right]\,.$
Note that the effect of neutrino asymmetries and spectral distortions is manifested by the neutrino density distribution term $\frac{{d}\ln\varrho_{i}^{0}}{{d}\ln q}$. Since the density matrices for neutrinos and antineutrinos with different flavors are different in the presence of primordial neutrino asymmetries, the Boltzmann hierarchies for the corresponding perturbations $\Psi_{i}$ and $\bar{\Psi}_{i}$ are in general different and should be treated separately. With the moments $\Psi_{i,l}$, the perturbations in physical quantities, such as the energy density, pressure, energy flux and shear stress, can be computed straightforwardly by integrating different moments. For our purpose, the fractional density perturbation for the neutrinos with family number $i$ can be expressed as
[…]
where $\rho_{i,0}(t)$ is the zeroth-order homogeneous background energy density for neutrinos. Eventually, the effects of neutrino masses and asymmetries at different length scales can be reflected in the evolution of the Fourier modes $\delta_{i,k}$, which we obtain via the CLASS code.
5.1.3 Effects on the BAO
Following Ma:1995ey , the evolution of the perturbations for photon obey the Boltzmann hierarchy,
$\dot{\delta}_{\gamma}$ $=$ $-{4\over 3}\theta_{\gamma}-{2\over 3}\dot{h}\,,$
$\dot{\theta}_{\gamma}$ $=$ $k^{2}\left(\frac{1}{4}\delta_{\gamma}-\sigma_{\gamma}\right)+an_{e}\sigma_{T}(\theta_{b}-\theta_{\gamma})\,,$
[…]
$\dot{F}_{\gamma\,l}$ $=$ ${k\over 2l+1}\left[lF_{\gamma\,(l-1)}-(l+1)F_{\gamma\,(l+1)}\right]-an_{e}\sigma_{T}F_{\gamma\,l}\,,\quad l\geq 3\,,$
[…]
where $\delta_{\gamma}$ is the density fluctuation of photon, $\theta_{\gamma}$ is the divergence of photon velocity, and ${F}_{\gamma\,l}$ is the density perturbation of photon. In addition, $n_{e}$ is the electron number density and $\sigma_{T}$ is the Thomson scattering cross section. The Boltzmann equations for baryon are
$\dot{\delta}_{b}$ $=$ $-\theta_{b}-{1\over 2}\dot{h}\,,$
$\dot{\theta}_{b}$ $=$ $-{\dot{a}\over a}\theta_{b}+c_{s}^{2}k^{2}\delta_{b}+{4\bar{\rho}_{\gamma}\over 3\bar{\rho}_{b}}an_{e}\sigma_{T}(\theta_{\gamma}-\theta_{b})\,,$ (100)
where $\delta_{b}$ and $\theta_{b}$ is the baryon density fluctuation and the divergence of baryon fluid velocity, and $c_{s}^{2}$ is the square of the baryon sound speed.
Due to the electromagnetic interactions between photons and electrons, they evolve differently before and after recombination. Before recombination, photons and baryons are tightly coupled so that they can be treated as a single photon-baryon fluid, called as the "tight-coupling approximation". During such "tight-coupling epoch", the perturbations of baryon and photon undergo harmonic motion with very slowly decaying amplitude (Peebles:1970ag, ; Sunyaev:1970bma, ). After recombination, photons release their hold on the baryons, so that the baryon acoustic oscillations (BAO) become frozen-in, and the the baryon perturbations begin to grow as $\propto a$. The BAO feature then gets transferred to the gravitational potential and the matter fluctuations, leaving imprints in the galaxy power spectrum that depends on the sound horizon at decoupling and the angular diameter distance to the surface of photon-electron last-scattering.
In the presence of primordial neutrino asymmetries, the matter-radiation equality will be directly modified by the additional contribution to $N_{\rm eff}$, leading to the modification of the BAO properties. In addition, the effects of primordial neutrino asymmetries on the evolution of neutrino perturbations are also transported to photons and baryons through metric perturbations encoded in $h$ and $\eta$, even though neutrinos are collisionless. As a consequence, the primordial neutrino asymmetries have significant implications for both the CMB angular power spectrum and the BAO in the matter power spectrum, as we will see later.
5.2 Results of the CMB anisotropies and the matter power spectrum
Figure 7: Left panel: The CMB angular power spectrum $C_{l}^{\rm TT}$ within the SM case for different total neutrino masses, compared to the results with massless neutrinos. Right panel: The linear matter power spectrum $P(k)$ within the SM case for different total neutrino masses, compared to the results with massless neutrinos.
We present in this subsection the numerical results of the CMB and LSS obtained via the CLASS code. Since both observable are affected by the neutrino masses and asymmetries at the same time, we first discuss the effect of neutrino masses on the CMB and LSS within the SM. We present the ratio of the CMB angular power spectrum between the results with massive neutrinos and massless neutrinos $C_{l}^{\rm TT}/C_{l}^{\rm TT}(\sum m_{\nu}=0)$ and the similar ratio for the matter power spectrum $P_{k}/P_{k}(\sum m_{\nu}=0)$ for several choices of total neutrino masses in the left and right panels of Fig. 7, respectively. Note that the normal order of the neutrino mass hierarchy is considered in this work, so we adjust the $\mu_{\nu_{1}}$ for each choice of total mass, while $\mu_{\nu_{2}}$ and $\mu_{\nu_{3}}$ are obtained by the fixed mass differences between neutrinos. Focusing on the left panel, the ratio of the $C_{l}^{\rm TT}$ between massive and massless neutrinos can be divided into three stages. At large angular scales (lower multipoles ${\ell\lesssim 20}$), massive neutrinos induce a small suppression which comes from the late time integrated Sachs-Wolfe effect. On the intermediate scales (${20\lesssim\ell\lesssim 500}$), dominated by the imprints of the acoustic oscillations in the baryon-photon fluid, the first peak of $C_{l}^{\rm TT}$ is affected by the modified evolution of the gravitational potential due to the correlation between massive neutrinos and metric perturbations. As we can see, the height of the first peak given by the early integrated Sachs-Wolfe effect is increased for larger neutrino masses compared to the case with massless neutrinos. The acoustic peaks at small scales (${\ell\gtrsim 500}$) are also affected by the massive neutrinos. Similar with the first peak, the height of the acoustic peaks is increased for larger neutrno masses.
Similarly, the resulting matter power spectrum with massive neutrinos can also be divided to three stages. At large scales corresponding to $k\lesssim 10^{-3}\,h/{\rm Mpc}$, neutrinos do not enter the horizon, and the small suppression of $P(k)$ is induced by the modification of the matter-radiation equality from the non-relativistic transition of massive neutrinos. At intermediate scales corresponding to $10^{-3}\,h/{\rm Mpc}\lesssim k\lesssim 1\,h/{\rm Mpc}$, neutrinos gradually enter the horizon, so that the free streaming of these neutrinos results in an additional suppression of the matter power spectrum. Therefore, $P_{k}/P_{k}(\sum m_{\nu}=0)$ decreases with increasing wavenumber $k$. And at small scales ($k\gtrsim 1\,h/{\rm Mpc}$), the majority of neutrinos stream freely, leading to an almost constant suppression of the power spectrum. The larger the total neutrino mass, the greater the suppression of the matter power spectrum. It is also worth noting that the matter power spectrum is more sensitive to the mass of the neutrinos, as can be seen from the plots. Therefore, it is more efficient to constrain the neutrino masses with the LSS observations, as we will see in the next section.
Figure 8: Left panel: The CMB angular power spectrum $C_{l}^{\rm TT}$ for different neutrino degeneracy parameters $\xi_{\nu}$, compared to the results with $\xi_{\nu}=0$. Right panel: The linear matter power spectrum $P(k)$ for different neutrino degeneracy parameters $\xi_{\nu}$, compared to the results with $\xi_{\nu}=0$. The total neutrino mass is fixed to $\sum m_{\nu}=0.1{\rm eV}$ for both panels.
Figure 9: The sound horizon at the baryon drag epoch of the early universe $r_{d}$ as a function of the primordial neutrino asymmetry $\xi_{\nu}$.
We now present the results for non-zero $\xi_{\nu}$. To isolate the effects of neutrino asymmetries from the total neutrino mass, the left panel of Fig. 8 shows the ratio of the CMB angular power spectrum for $\xi_{\nu}\neq 0$ to $\xi_{\nu}=0$, with $\sum m_{\nu_{i}}=0.1\,\text{eV}$. A similar ratio for the matter power spectrum is shown in the right panel. As seen in the left panel, the effects of neutrino asymmetries on the CMB angular power spectrum manifest primarily at small scales ($\ell\gtrsim 400$). At $\ell\lesssim 400$, $C_{l}^{\rm TT}$ for non-zero $\xi_{\nu}$ matches the results for $\xi_{\nu}=0$. However, as $\ell$ increases, deviations due to $\xi_{\nu}$ become apparent. These deviations depend on both the magnitude and sign of $\xi_{\nu}$: positive $\xi_{\nu}$ enhances the small-scale power, while negative $\xi_{\nu}$ leads to suppression. This behavior is largely driven by changes in helium abundance, which alters the CMB spectrum via diffusion damping. In addition to overall enhancement or suppression, the acoustic peaks are also affected by $\xi_{\nu}$, as larger $\xi_{\nu}$ values modify the oscillations at the last scattering surface. The amplitude of these modifications grows with the absolute value of $\xi_{\nu}$, with negative $\xi_{\nu}$ producing more pronounced effects.
For the linear matter power spectrum in the right panel of Fig. 8, we observe that at large scales, the result for non-zero $\xi_{\nu}$ is identical to that for $\xi_{\nu}=0$. At small scales, however, the free-streaming of neutrinos is enhanced with larger $\xi_{\nu}$, resulting in a slightly stronger suppression of the matter power spectrum, although this suppression is insensitive to the sign of $\xi_{\nu}$. The more intriguing effect of non-zero $\xi_{\nu}$ is the oscillations of the matter power spectrum on the scales $10^{-1}\,h/{\rm Mpc}\gtrsim k\gtrsim 10^{0}\,h/{\rm Mpc}$, which represents the effect of the BAO. Similar to the acoustic oscillations in the CMB angular power spectrum, we find that the amplitudes of the BAO in the matter power spectrum are also larger with a larger absolute value of $\xi_{\nu}$, while the negative $\xi_{\nu}$ leads to a larger effect than that for the positive $\xi_{\nu}$. In addition to the imprints of BAO on the matter power spectrum, we show in Fig. 9 the sound horizon at the baryon drag epoch of the early universe $r_{d}$ as a function of the primordial neutrino asymmetry $\xi_{\nu}$. Due to the additional contribution to $N_{\rm eff}$ for non-zero $\xi_{\nu}$ according to Eq. 63, the larger the $\left|\xi_{\nu}\right|$, the smaller the $r_{d}$, resulting in the significant imprints on the matter power spectrum. These results imply that BAO experiments can also provide powerful constraints on the primordial neutrino asymmetry, as we will in Section 6, which is often overlooked in previous studies. Note also that these conclusions are drawn with a fixed total neutrino mass $\sum m_{\nu_{i}}=0.1\,\text{eV}$. However, they remain valid for different total neutrino masses, although the resulting ratios $C_{l}^{\rm TT}/C_{l}^{\rm TT,\,\xi_{\nu}=0}$ and $P(k)/P^{\xi_{\nu}=0}(k)$ may vary slightly at intermediate scales due to the non-relativistic transition of massive neutrinos.
Figure 10: The CMB angular power spectrum $C_{l}^{\rm TT}$ (left panel) and the linear matter power spectrum $P(k)$ (right panel) for different neutrino degeneracy parameters $\xi_{\nu}$, compared to the corresponding results with $\xi_{\nu}=0$. The total neutrino mass is fixed to $\sum m_{\nu}=0.1{\rm eV}$ for both panels. The solid curves are the same with Fig. 5, corresponding to the results including the spectral distortions of the neutrinos and antineutrinos, while the dashed curves correspond to the results without the spectral distortions.
Finally, we briefly discuss the effects of spectral distortions of the neutrinos and antineutrinos on the CMB and LSS. Similar to Fig. 8, we present the resulting CMB angular power spectrum $C_{l}^{\rm TT}$ and linear matter power spectrum $P(k)$ with non-zero $\xi_{\nu}$, neglecting the neutrino spectral distortions, as the dashed curves in the left and right panels of Fig. 10. The solid curves in Fig. 10 are the same as those in Fig. 8, corresponding to the results including the spectral distortions of the neutrinos and antineutrinos. Comparing the resulting CMB angular power spectrum $C_{l}^{\rm TT}$ with and without the neutrino spectral distortions, we see that the primary effect of including the spectral distortions is a slight suppression of the first three acoustic peaks, with an almost negligible effect on the remaining peaks in the small scales. For the linear power spectrum, we find that the effect of spectral distortions on the BAO is similar with that on the CMB angular power spectrum. Here the first two acoustic peaks are slightly enhanced while the effects on the remaining peaks are negligible. In addition, we find that the calculations without the neutrino spectral distortions cannot capture the right behavior of the matter power spectrum at small scales, where the results without the spectral distortions predict a near constant suppression at the scales $k\gtrsim 1\,h/{\rm Mpc}$, while the preciser results with the spectral distortions show that the matter power spectrum receives larger suppression at these scales. However, the discussion here refers only to the linear matter power spectrum, and we expect that a comprehensive study of the effects of neutrino spectral distortions on the LSS in the nonlinear regime can be done in the future with the N-body simulations.
6 Combined Analysis of Primordial Neutrino Asymmetries
In previous sections, we have shown that the primordial neutrino asymmetry significantly influences the neutrino decoupling process, leaving significant observable imprints on the Universe, such as the abundance of light elements produced during the BBN, the CMB anisotropies, and the BAO imprints in the LSS. These cosmological observations can then be used to constrain the primordial neutrino asymmetries. In this section, we consolidate the results of the previous sections and discuss the constraints on the primordial neutrino asymmetry derived from contemporary observations of the BBN, CMB and BAO. In addition, we briefly explore the corresponding constraints on UV model parameters using a specific model that generates primordial neutrino asymmetries in the early universe.
6.1 Observational Constraints on Primordial Neutrino Asymmetry
Bounds for the model $\mathbf{\Lambda}$CDM + $\mathbf{\xi_{\nu}}$ + $\mathbf{m_{\nu_{1}}}$
Data Sets $\qquad\,\,\,{\bf\xi_{\nu}}\,\,\,\qquad$ $\omega_{b}$ $\omega_{cdm}$ ${\sum m_{\nu}\,\,(\mathrm{eV})}$
BBN $0.039\pm 0.013$ $2.155\pm 0.016$
BBN + CMB $0.028\pm 0.012$ $2.189\pm 0.012$ $0.1225\pm 0.0013$ 0.414
BBN + CMB + BAO $0.024\pm 0.012$ $2.200\pm 0.011$ $0.1205\pm 0.0009$ 0.131
Bounds from the datasets BBN + CMB + BAO
Models $\qquad\,\,\,{\bf\xi_{\nu}}\,\,\,\qquad$ $\omega_{b}$ $\omega_{cdm}$ ${\sum m_{\nu}\,\,(\mathrm{eV})}$
$\mathbf{\Lambda}$CDM + $\mathbf{m_{\nu_{1}}}$ $2.211\pm 0.010$ $0.1203\pm 0.0009$ 0.136
$\mathbf{\Lambda}$CDM + $\mathbf{\xi_{\nu}}$ $0.023\pm 0.012$ $2.200\pm 0.011$ $0.1207\pm 0.0009$
$\mathbf{\Lambda}$CDM + $\mathbf{\xi_{\nu}}$ + $\mathbf{m_{\nu_{1}}}$ $0.024\pm 0.012$ $2.200\pm 0.011$ $0.1205\pm 0.0009$ 0.131
Table 3: Summary of constraints on the primordial lepton asymmetry $\xi_{\nu}$, the baryon abundance $\omega_{b}$, the cold dark matter abundance $\omega_{cdm}$ and the total neutrino mass $\sum m_{\nu}$ from various datasets and cosmological models.
In this work we utilize the following datasets for BBN, CMB, and BAO in our analysis:
BBN
The EMPRESS measurement of helium abundance, ${Y\rm_{P}}=0.2370^{+0.0034}_{-0.0033}$ empress , and the deuterium abundance adopted by the PDG group, $10^{5}\times{\rm D/H}=2.547\pm 0.025$ pdg , which is based on the results of Cooke:2017cwo .
CMB
The Planck 2018 low-$l$ and high-$l$ TT, TE, and EE spectra, along with the reconstructed CMB lensing spectrum Planck:2018vyg ; Planck:2018lbu ; Planck:2019nip .
BAO
The BOSS DR12 BOSS:2016wmc with its full covariance matrix for BAO, as well as the 6dFGS Beutler:2011hx and MGS of SDSS Ross:2014qpa .
We note that the helium abundance from the BBN data set primarily constrains the primordial neutrino asymmetries, while the deuterium abundance is more sensitive to the baryon density $\omega_{b}$. The BAO dataset is mainly used to constrain the total neutrino mass $\sum m_{\nu}$, but as we will discuss, it also places additional constraints on primordial neutrino asymmetries.
To extract the constraints from these datasets, we conduct MCMC analyses using the Montepython engine Audren:2012wb ; Brinckmann:2018cvx . For analysis, the $N_{\rm eff}$ is adopted as Eq. (63), derived in Section 3. The helium and deuterium abundances predicted by BBN are computed using our fitting formulas in Equation (85). As shown in Fig. 10, the effects of neutrino spectral distortions on the CMB angular power spectrum and the BAO in the matter power spectrum are small. Therefore, for simplicity, the (anti)neutrino spectral distortions are neglected in the MCMC analysis. The parameter set for the CMB and BAO datasets includes the six independent parameters of the $\Lambda$CDM model, along with the neutrino degeneracy parameter $\xi_{\nu}$ and the mass of the lightest neutrino for the normal hierarchy $m_{\nu_{1}}$, i.e., $\{\omega_{b},\omega_{cdm},\theta_{s},A_{s},n_{s},\tau_{\mathrm{reio}},\xi_{\nu},m_{\nu_{1}}\}$, where $\theta_{s}$ represents the sound horizon, $A_{s}$ is the amplitude of the primordial curvature perturbations at the pivot scale $k=0.05\,\mathrm{Mpc}^{-1}$, $n_{s}$ is the scalar spectral index, and $\tau_{\mathrm{reio}}$ is the optical depth to reionization. On the other hand, we only consider $\{\omega_{b},\xi_{\nu}\}$ for the BBN dataset, as the predicted light element abundances from BBN are largely independent of other cosmological parameters within the relevant range.
Figure 11: 2D and 1D marginalised posterior distributions for the baryon abundance $\omega_{b}$, the neutrino degeneracy parameter $\xi_{\nu}$, the cold dark matter abundance $\omega_{cdm}$, the mass of the lightest neutrino for the normal hierarchy $m_{\nu_{1}}$, and scalar spectral index $n_{s}$. The red, blue, and green regions represent constraints from BBN, CMB + BAO, and their combination, respectively.
Figure 12: 1D marginalised distributions for different parameter sets derived from the BBN + CMB + BAO datasets.
With the aforementioned choices of parameters and datasets, we modified the MontePython sampler and CLASS codes to perform the MCMC analysis. The Gelman-Rubin criterion for all chains converged to $R-1<0.01$. The resulting $1\,\sigma$ cosmological constraints are summarized in Table 3, where the $95\%$ confidence limits on the total neutrino mas are also presented from the constraints on the mass of the lightest neutrino for the normal hierarchy $m_{\nu_{1}}$. For the same model $\Lambda$CDM + ${\xi_{\nu}}$ + ${m_{\nu_{1}}}$, we present the 1D and 2D marginalized posterior distributions of key parameters that show significant variations across different datasets in Fig. 11, including the baryon abundance $\omega_{b}$, the neutrino degeneracy parameter $\xi_{\nu}$, the cold dark matter abundance $\omega_{cdm}$, the sound horizon $\theta_{s}$, and the mass of the lightest neutrino for the normal hierarchy $m_{\nu_{1}}$.
Concerning the constraints on the primordial neutrino asymmetries, we see that the recent EMPRESS data in the BBN dataset points to a positive primordial neutrino asymmetry $\xi_{\nu}=0.039\pm 0.013$, which is different from zero with a $3.0\,\sigma$ significance. Combining the BBN and CMB data, the resulting constraints on $\xi_{\nu}$ are significantly reduced to $\xi_{\nu}=0.028\pm 0.012$, different from zero with a $2.3\,\sigma$ significance. Compared to the constraints $\xi_{\nu}=0.034\pm 0.014$ reported in previous study with the same BBN and CMB datasets Escudero:2022okz , our result indicate a $\sim 18\%$ smaller central value and a slight reduction in the uncertainty. Incorporating BOSS BAO observations in the MCMC analysis, the primordial neutrino asymmetry constraint further reduce to $\xi_{\nu}=0.024\pm 0.013$, indicating a $\sim 29\%$ reduction in the central value compared to previous results, while the significance of a positive neutrino asymmetry is reduced to $2.0\,\sigma$. Our study demonstrate the additional constraining power of the BAO experiments on the primordial neutrino asymmetry, which are not aware in previous studies.
We find that the constraints on other parameters also vary significantly with different datasets, especially for $\omega_{b}$, $\omega_{cmb}$, $n_{s}$ and $m_{\nu_{1}}$, as shown in Fig. 11. For the constraints on $\omega_{b}$, the difference is largely due to the fact that we chose the nuclear reaction rates in PRIMAT in the process of BBN, which predicted a deuterium abundance lower than the current constraints from PDG, thus leading to significantly lower constraints on $\omega_{b}$ Pitrou:2018cgg ; Burns:2022hkq ; Escudero:2022okz . While the constraints on the primordial neutrino asymmetry $\xi_{\nu}$ do not strongly depend on the choice of nuclear reaction rates, as discussed in  Escudero:2022okz , a detailed study with different nuclear reaction rates will be performed in our companion paper. The inclusion of the BOSS BAO data significantly changes the corresponding constraints on $\omega_{cmb}$, $n_{s}$ and $m_{\nu_{1}}$ due to its sensitivity to the evolution of the late universe. Especially, the resulting $95\%$ upper limit for the total mass of neutrinos are greatly tightened from $0.414$ eV by BBN + CMB to $0.131$ eV by BBN + CMB + BAO, since the BAO data help to break the CMB degeneracy. Recently, the BAO observations of DESI have put an even stronger bound on neutrino masses $\sum m_{\nu_{i}}\lesssim 0.07\,eV$ DESI:2024mwx . The relevant studies on primordial neutrino asymmetry and neutrino masses with DESI BAO data will also be discussed in our companion paper.
Finally we briefly discuss the effects of different parameter sets on the resulting constraints, especially the interplay between the primordial neutrino asymmetry and the total mass of neutrinos. Using the data set of BBN + CMB + BAO, we present in Fig. 12 the resulting 1d marginalised distributions for three choices of the parameter set, first is the set of $\Lambda CDM+m_{\nu_{1}}$, second is the $\Lambda CDM+\xi_{\nu}$, while the third is $\Lambda CDM+m_{\nu_{1}}+\xi_{\nu}$. As we can see, while the different choices of the parameter set lead to quite modifications to the constraints on the six parameters of $\Lambda CDM$, especially for $\omega_{b}$ and $n_{s}$, the resulting constraints on the additional parameters are almost the same for different parameter sets. The constraints on the total mass of neutrinos are only slightly tightened while including primordial neutrino asymmetry $\xi_{\nu}$ as a free parameter, while the constraints on $\xi_{\nu}$ are almost the mass as the inclusion of free $\sum m_{\nu}$. Therefore, we conclude that the constraints on neutrino masses and the primordial neutrino asymmetries can indeed be analysed separately if one do not concern the constraints on the other parameters in $\Lambda CDM$, otherwise the parameter sets with both the neutrino masses and the primordial neutrino asymmetries must be considered.
6.2 Constraints on UV model parameters
With the constraints on the primordial neutrino asymmetries, we discuss in this subsection the corresponding constraints on the UV model parameters with a specific model to generate the primordial neutrino asymmetries. Here we consider a Q-ball decay model Kawasaki:2022hvx , where Q-balls are produced in the Affleck-Dine mechanism Kusenko:1997si ; Enqvist:1997si ; Kasuya:1999wu , which utilizes flat directions in the minimal supersymmetric standard model (MSSM). When some flat direction with lepton number has a large field value during inflation, it generates lepton number by field dynamics after inflation. At the same time, Q-balls are produced and the generated lepton number is all confined inside the Q-balls. Thus, such Q-balls have a large lepton number and are called L-balls. Since the sphaleron effect is inactive inside the L-balls, the lepton asymmetry produced by the Affleck-Dine mechanism are kept without being converted to a baryon asymmetry Kawasaki:2002hq ; Gelmini:2020ekg . The L-balls decay into neutrinos below the electroweak scale and realize a large lepton asymmetry in the universe.
Following the results of Kawasaki:2022hvx , the primordial lepton asymmetry $\eta_{L}$ is defined as the ratio between the difference of the number density of neutrinos over antineutrinos and the photon number density,
$\eta_{L}\equiv\sum_{\alpha}\frac{n_{\nu_{\alpha}}-n_{\bar{\nu}_{\alpha}}}{s}\simeq 0.105\,\xi_{\nu},$ (101)
where $s=\frac{2\pi^{2}}{45}g_{\star}T_{\gamma}^{3}$ is the entropy density and $g_{\star}=10.75$ is the relativistic degree of freedom at the neutrino decoupling epoch. In addition, the resulting $\eta_{L}$ in the Q-ball decay model can be expressed as
$\eta_{L}=\frac{3T_{\mathrm{D}}}{4m_{3/2}},$ (102)
where $m_{3/2}$ is the the gravitino mass, $T_{\mathrm{D}}$ is the decay temperature of the L-balls and can be expressed with the other UV model parameters with
$T_{\mathrm{D}}$ $\simeq 2.69~{}\mathrm{MeV}\,$
×$\hskip 15.0pt\times\left(\frac{g_{*}}{10.75}\right)^{-1/4}\left(\frac{\beta}{6\times 10^{-4}}\right)^{-5/8}\left(\frac{m_{3/2}}{0.5~{}\mathrm{GeV}}\right)^{5/2}\left(\frac{M_{F}}{5\times 10^{6}~{}\mathrm{GeV}}\right)^{-2}\left(\frac{N_{\ell}}{3}\right)^{1/2}\left(\frac{\zeta}{2.5}\right)^{1/2},$ (103)
where $g_{*}(T_{\mathrm{D}})$ is the relativistic number of degrees of freedom at $T_{\mathrm{D}}$, $\beta\simeq 6\times 10^{-4}$ is a dimensionless constant, $M_{F}$ is the SUSY breaking scale, $N_{\ell}$ is the number of decay channels. Therefore, once we have obtained the constraints on the lepton asymmetry from cosmological observations, they can be directly translated into constraints on the model parameters of the UV model.
According to our study, the constraints on $\eta_{L}$ from the BBN + CMB + BAO datasets are $\eta_{L}=0.0025\pm 0.0013$. Therefore, the constraint on the L-balls decay temperature is
$T_{\mathrm{D}}$ $=\frac{4}{3}\eta_{L}m_{3/2}\simeq 3.3~{}\mathrm{MeV}\,\left(\frac{m_{3/2}}{0.5~{}\mathrm{GeV}}\right)\left(\frac{\eta_{L}}{5\times 10^{-3}}\right)$
[…]
and conversely,
$m_{3/2}$ $=\frac{3T_{\mathrm{D}}}{4T\eta_{L}}\simeq 0.5~{}\mathrm{GeV}\,\left(\frac{T_{\mathrm{D}}}{3.3~{}\mathrm{MeV}}\right)\left(\frac{5\times 10^{-3}}{\eta_{L}}\right)$
[…]
These constraints can also be converted into the constraints on other UV model parameters, such as the SUSY breaking scale, as seen in the preliminary study of Kawasaki:2022hvx .
7 Conclusion
The presence of primordial neutrino asymmetries before the neutrino decoupling epoch has significant implications for both the neutrino decoupling process and subsequent cosmological phenomena, and can therefore be constrained by observations of BBN, CMB and LSS. However, the effects of primordial neutrino asymmetries on neutrino decoupling—particularly concerning the resulting $N_{\rm eff}$ and the spectral distortions of (anti)neutrinos, including the impacts of flavor oscillations and FTQED corrections—remain ambiguous. These ambiguities affect the interpretation of subsequent processes. Consequently, studies of primordial neutrino asymmetries in relation to BBN and the CMB often employ approximations, such as using thermal Fermi-Dirac distributions for (anti)neutrinos, without thorough scrutiny.
In this work, we aim to address these ambiguities by numerically solving the neutrino decoupling process and analyzing the implications of the resulting (anti)neutrino density matrices for BBN, CMB, and LSS. We begin by re-deriving the quantum kinetic equations (QKEs) for neutrinos and antineutrinos using the Closed Time Path (CTP) formalism for two-point functions. Focusing on primordial neutrino asymmetries with complete flavor equilibration, where $\xi_{\nu_{e}}=\xi_{\nu_{\mu}}=\xi_{\nu_{\tau}}=\xi_{\nu}$, we numerically solve the QKEs in the early universe using the public code FortEPiaNO Gariazzo:2019gyi ; Bennett:2020zkv , including the effects of neutrino oscillations, matter effects, FTQED corrections up to $\mathcal{O}(3)$, and comprehensive neutrino-electron and neutrino-neutrino collision terms. The final $N_{\rm eff}$ is then evaluated based on the resulting density matrices, and can be expressed as $N_{\rm eff}\simeq N_{\rm eff}^{\rm SM}+3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right)+0.0102\xi_{\nu}^{2}$, where $N_{\rm eff}^{\rm SM}=3.0440\pm 0.0002$, while the second and last term represent the corrections from nonzero $\xi_{\nu}$ in the instantaneous decoupling limit and from the non-instantaneous effect, respectively. We also confirm that the spectral distortions of neutrinos and antineutrinos with non-zero $\xi_{\nu}$ differ due to distinct Pauli-blocking effects, in agreement with the priliminary study in Grohs:2016cuu , where neutrino spectral distortions were evaluated while neglecting flavor oscillations and FTQED corrections.
Using the density matrices for neutrinos and antineutrinos, we further investigate the implications of primordial neutrino asymmetries for BBN, CMB and LSS. For BBN, we introduce a novel correction to the weak interaction rates arising from the distinct spectral distortions of neutrinos and antineutrinos. The light element abundances are then computed with the public BBN code PRIMAT Pitrou:2018cgg , incorporating the other additional corrections on weak rates. According to the the EMPRESS measurement on the helium abundance $Y_{P}|_{\rm EMPRESS}=0.2370^{+0.0034}_{-0.0033}$, our results immediately indicate a positive primordial neutrino asymmetry $0.032\leq\xi_{\nu}\leq 0.052$ with fixed baryon abundance $\omega_{b}$. The effects of the new proposed correction are also examined, and we find that the spectral distortions for neutrinos and antineutrinos are important for precise predictions for BBN, especially for the abundance of Helium-4. Compared with the results in the instantaneous limit, the traditional treatment with Fermi-Dirac distributions predict a negative relative difference for a positive primordial neutrino asymmetry, while our precise treatment with spectral distortions show that the relative difference is still positive. Additionally, we have provided semi-analytic expressions for these abundances as functions of $\xi_{\nu}$, which will be useful for future studies. For the CMB and LSS, a novel finding of our study is that primordial neutrino asymmetries also affect the amplitudes of the acoustic oscillations, both for the CMB and for baryon acoustic oscillations (BAO). This shows that BAO observations could serve as a tool for constraining primordial neutrino asymmetries.
Based on our systematic understanding of primordial asymmetries, we perform a joint MCMC analysis including the neutrino degeneracy parameter $\xi_{\nu}$ and the total neutrino mass $\sum m_{\nu}$, using data from EMPRESS BBN, Planck CMB, and BOSS BAO observations. Combining the EMPRESS BBN and Planck CMB datasets, the primordial neutrino asymmetry constraint yields $\xi_{\nu}=0.028\pm 0.013$, indicating a $\sim 18\%$ smaller central value and a slight reduction in the uncertainty compared to previous results $\xi_{\nu}=0.034\pm 0.014$ obtained in Escudero:2022okz . Incorporating BOSS BAO observations in the MCMC analysis, the primordial neutrino asymmetry constraint further reduce to $\xi_{\nu}=0.024\pm 0.013$, indicating a $\sim 29\%$ reduction in the central value compared to previous results and a $2.0\,\sigma$ significance of a positive neutrino asymmetry. We also discuss the impacts of treating the sum of neutrino mass $\sum m_{\nu}$ and the primordial neutrino asymmetry $\xi_{\nu}$ as free parameters on the resulting constraints for each of them. Finally, we also show how the constraints on the neutrino asymmetries can be translated into constraints on the UV model parameters with a specific model in which the large neutrino asymmetries are generated by the decay of Q-balls Kawasaki:2002hq ; Gelmini:2020ekg .
Our work provides the framework to obtain the implications of new physics beyond the SM and $\Lambda$CDM model for neutrino decoupling and subsequent processes, and the corresponding cosmological constraints from present and future BBN, CMB and LSS observations. Thus, our work facilitates a more comprehensive analysis for primordial neutrino asymmetry and other models with new physics, such as precise consideration of the neutrino synchronous oscillations Froustey:2021azz ; Froustey:2024mgf or neutrino non-standard interactions Du:2021idh ; Du:2023upj , based on more and more precise cosmological observations in the future.
Acknowledgements.
This work is supported by the National Science Foundation of China under Grants No. 12347105, No. 12375099 and No. 12047503, and the National Key Research and Development Program of China Grant No. 2020YFC2201501, No. 2021YFA0718304. The authors gratefully acknowledge the use of publicly available codes FortEPiaNO Gariazzo:2019gyi ; Bennett:2020zkv , PRIMAT Pitrou:2018cgg , CLASS Lesgourgues:2011re ; Lesgourgues:2011rh and MontePythonAudren:2012wb ; Brinckmann:2018cvx .
Appendix A Collision Terms for Neutrinos and Antineutrinos
A.1 Complete expressions for the collision terms
In this work, we consider the interaction processes of neutrino-electron/positron scattering ($\nu{}e\leftrightarrow\nu{}e$) and annihilation ($\nu\bar{\nu}\leftrightarrow e^{+}e^{-}$), plus neutrino-neutrino self-interactions ($\nu\nu\leftrightarrow\nu\nu$, $\nu\bar{\nu}\leftrightarrow\nu\bar{\nu}$). Therefore, the total collision term for neutrinos $\mathcal{I}$ can be decomposed as
$\mathcal{I}\equiv\mathcal{I}_{\rm sc}+\mathcal{I}_{\rm ann}+\mathcal{I}_{\nu\nu}$ (106)
where $\mathcal{I}_{\rm sc}$ and $\mathcal{I}_{\rm ann}$ represent the collision terms for neutrino-electron/positron scattering and annihilation respectively, while $\mathcal{I}_{\nu\nu}$ represents the collision terms for neutrino self-interactions. The collision term for antineutrinos $\bar{\mathcal{I}}$ also admits a similar decomposition. And the contributions from other interaction processes such as $\mu^{\pm}$ scattering and annihilation are neglected, as their contributions are highly suppressed at the temperature of neutrino decoupling.
In the following we present the explicit form of the collision terms, which are also presented in various studies (see $e.g.,$  Blaschke:2016xxt ). The collision terms for the neutrino-electron/positron scattering processes ($\nu{}e\leftrightarrow\nu{}e$) and the annihilation processes ($\nu\bar{\nu}\leftrightarrow e^{+}e^{-}$) take the form as
$\mathcal{I}_{\rm sc}=$ $\frac{1}{2}\frac{2^{5}G_{F}^{2}}{2p_{1}}\int{\mathrm{d}\pi_{2}\,\mathrm{d}\pi_{3}\,\mathrm{d}\pi_{4}\,(2\pi)^{4}\delta^{(4)}(p_{1}+p_{2}-p_{3}-p_{4})}$
$\Big{[}4(p_{1}\cdot p_{2})(p_{3}\cdot p_{4})\left(F_{\mathrm{sc}}^{LL}(\nu^{(1)},e^{(2)},\nu^{(3)},e^{(4)})+F_{\mathrm{sc}}^{RR}(\nu^{(1)},\bar{e}^{(2)},\nu^{(3)},\bar{e}^{(4)})\right)$
$+4(p_{1}\cdot p_{4})(p_{2}\cdot p_{3})\left(F_{\mathrm{sc}}^{RR}(\nu^{(1)},e^{(2)},\nu^{(3)},e^{(4)})+F_{\mathrm{sc}}^{LL}(\nu^{(1)},\bar{e}^{(2)},\nu^{(3)},\bar{e}^{(4)})\right)$
$-2(p_{1}\cdot p_{3})m_{e}^{2}\left(F_{\mathrm{sc}}^{LR}(\nu^{(1)},e^{(2)},\nu^{(3)},e^{(4)})+F_{\mathrm{sc}}^{LR}(\nu^{(1)},\bar{e}^{(2)},\nu^{(3)},\bar{e}^{(4)})+\{L\leftrightarrow R\}\right)\Big{]}\,,$ (107)
$\mathcal{I}_{\rm ann}=$ $\frac{1}{2}\frac{2^{5}G_{F}^{2}}{2p_{1}}\int{\mathrm{d}\pi_{2}\,\mathrm{d}\pi_{3}\,\mathrm{d}\pi_{4}\,(2\pi)^{4}\delta^{(4)}(p_{1}+p_{2}-p_{3}-p_{4})}$
$\Big{[}4(p_{1}\cdot p_{4})(p_{2}\cdot p_{3})F_{\mathrm{ann}}^{LL}(\nu^{(1)},\bar{\nu}^{(2)},e^{(3)},\bar{e}^{(4)})$
$+4(p_{1}\cdot p_{3})(p_{2}\cdot p_{4})F_{\mathrm{ann}}^{RR}(\nu^{(1)},\bar{\nu}^{(2)},e^{(3)},\bar{e}^{(4)})$
$+2(p_{1}\cdot p_{2})m_{e}^{2}\left(F_{\mathrm{ann}}^{LR}(\nu^{(1)},\bar{\nu}^{(2)},e^{(3)},\bar{e}^{(4)})+F_{\mathrm{ann}}^{RL}(\nu^{(1)},\bar{\nu}^{(2)},e^{(3)},\bar{e}^{(4)})\right)\Big{]}\,.$ (108)
where $\mathrm{d}\pi_{i}=\frac{\\ dd^{3}p_{i}}{(2\pi)^{3}2E_{i}}$ with $E_{i}=p_{i}$ if the particle $i$ is an (anti)neutrino or $E_{i}=\sqrt{p_{i}^{2}+m_{e}^{2}}$ if the particle $i$ is an electron/positron. The statistical factors incorporating the density distributions for all participating particles are
$F_{\mathrm{sc}}^{AB}(\nu^{(1)},e^{(2)},\nu^{(3)},e^{(4)})$ $=f_{e}^{(4)}(1-f_{e}^{(2)})\left[G^{A}\varrho^{(3)}G^{B}(1-\varrho^{(1)})+(1-\varrho^{(1)})G^{B}\varrho^{(3)}G^{A}\right]$
$-f_{e}^{(2)}(1-f_{e}^{(4)})\left[\varrho^{(1)}G^{B}(1-\varrho^{(3)})G^{A}+G^{A}(1-\varrho^{(3)})G^{B}\varrho^{(1)}\right]\,,$ (109)
$F_{\mathrm{ann}}^{AB}(\nu^{(1)},\bar{\nu}^{(2)},e^{(3)},\bar{e}^{(4)})$ $=f_{e}^{(3)}\bar{f}_{e}^{(4)}\left[G^{A}(1-\bar{\varrho}^{(2)})G^{B}(1-\varrho^{(1)})+(1-\varrho^{(1)})G^{B}(1-\bar{\varrho}^{(2)})G^{A}\right]$
$-(1-f_{e}^{(3)})(1-\bar{f}_{e}^{(4)})\left[G^{A}\bar{\varrho}^{(2)}G^{B}\varrho^{(1)}+\varrho^{(1)}G^{B}\bar{\varrho}^{(2)}G^{A}\right]\,,$ (110)
where $a,b\in\{L,R\}$ denote the charalitics for the electrons/positrons involved in the processes, $f(\bar{f})_{e}^{(i)}=1/(\exp{p_{i}/T_{\gamma}}+1)$ are the distribution functions for electrons/positrons, and $\varrho^{(i)},\bar{\varrho}^{(i)}$ are the corresponding density matrices for neutrinos and antineutrinos with momentum $p_{i}$. The interaction matrices used in the collision terms are
$G^{L}=\text{diag}(g_{L},\tilde{g}_{L},\tilde{g}_{L})\,,\qquad G^{R}=\text{diag}(g_{R},g_{R},g_{R})\,,$ (111)
where $g_{L}=\sin^{2}\theta_{W}+1/2$, $\tilde{g}_{L}=\sin^{2}\theta_{W}-1/2$, $g_{R}=\sin^{2}\theta_{W}$, and $\theta_{W}$ is the weak mixing angle.
The collision term for the neutrino self-interactions, including the scattering between neutrinos and (anti)neutrinos ($\nu\nu\leftrightarrow\nu\nu$, $\nu\bar{\nu}\leftrightarrow\nu\bar{\nu}$) and the annihilation between a pair of neutrinos and antineutrinos ($\nu\bar{\nu}\leftrightarrow\nu\bar{\nu}$), can be written as
$\mathcal{I}_{\nu\nu}=$ $\frac{1}{2}\frac{2^{5}G_{F}^{2}}{2p_{1}}\int{\mathrm{d}\pi_{2}\,\mathrm{d}\pi_{3}\,\mathrm{d}\pi_{4}\,(2\pi)^{4}\delta^{(4)}(p_{1}+p_{2}-p_{3}-p_{4})}$ (112)
$\Big{[}(p_{1}\cdot p_{2})(p_{3}\cdot p_{4})F_{\mathrm{sc}}(\nu^{(1)},\nu^{(2)},\nu^{(3)},\nu^{(4)})$
$+(p_{1}\cdot p_{4})(p_{2}\cdot p_{3})\left(F_{\mathrm{sc}}(\nu^{(1)},\bar{\nu}^{(2)},\nu^{(3)},\bar{\nu}^{(4)})+F_{\mathrm{ann}}(\nu^{(1)},\bar{\nu}^{(2)},\nu^{(3)},\bar{\nu}^{(4)})\right)\Big{]}\,,$
with the statistical factors for scattering and annihilation processes:
$F_{\mathrm{sc}}(\nu^{(1)},\nu^{(2)},\nu^{(3)},\nu^{(4)})=\\ \left[\varrho^{(4)}(\mathbb{1}-\varrho^{(2)})+\mathrm{Tr}(\cdots)\right]\varrho^{(3)}(\mathbb{1}-\varrho^{(1)})+(\mathbb{1}-\varrho^{(1)})\varrho^{(3)}\left[(\mathbb{1}-\varrho^{(2)})\varrho^{(4)}+\mathrm{Tr}(\cdots)\right]\\ -\left[(\mathbb{1}-\varrho^{(4)})\varrho^{(2)}+\mathrm{Tr}(\cdots)\right](\mathbb{1}-\varrho^{(3)})\varrho^{(1)}-\varrho^{(1)}(\mathbb{1}-\varrho^{(3)})\left[\varrho^{(2)}(\mathbb{1}-\varrho^{(4)})+\mathrm{Tr}(\cdots)\right]\,,$ (113)
$F_{\mathrm{sc}}(\nu^{(1)},\bar{\nu}^{(2)},\nu^{(3)},\bar{\nu}^{(4)})=\\ \left[(\mathbb{1}-\bar{\varrho}^{(2)})\bar{\varrho}^{(4)}+\mathrm{Tr}(\cdots)\right]\varrho^{(3)}(\mathbb{1}-\varrho^{(1)})+(\mathbb{1}-\varrho^{(1)})\varrho^{(3)}\left[\bar{\varrho}^{(4)}(\mathbb{1}-\bar{\varrho}^{(2)})+\mathrm{Tr}(\cdots)\right]\\ -\left[\bar{\varrho}^{(2)}(\mathbb{1}-\bar{\varrho}^{(4)})+\mathrm{Tr}(\cdots)\right](\mathbb{1}-\varrho^{(3)})\varrho^{(1)}-\varrho^{(1)}(\mathbb{1}-\varrho^{(3)})\left[(\mathbb{1}-\bar{\varrho}^{(4)})\bar{\varrho}^{(2)}+\mathrm{Tr}(\cdots)\right]\,,$ (114)
$F_{\mathrm{ann}}(\nu^{(1)},\bar{\nu}^{(2)},\nu^{(3)},\bar{\nu}^{(4)})=\\ \left[\varrho^{(3)}\bar{\varrho}^{(4)}+\mathrm{Tr}(\cdots)\right](\mathbb{1}-\bar{\varrho}^{(2)})(\mathbb{1}-\varrho^{(1)})+(\mathbb{1}-\varrho^{(1)})(\mathbb{1}-\bar{\varrho}^{(2)})\left[\bar{\varrho}^{(4)}\varrho^{(3)}+\mathrm{Tr}(\cdots)\right]\\ -\left[(\mathbb{1}-\varrho^{(3)})(\mathbb{1}-\bar{\varrho}^{(4)})+\mathrm{Tr}(\cdots)\right]\bar{\varrho}^{(2)}\varrho^{(1)}-\varrho^{(1)}\bar{\varrho}^{(2)}\left[(\mathbb{1}-\bar{\varrho}^{(4)})(\mathbb{1}-\varrho^{(3)})+\mathrm{Tr}(\cdots)\right]\,,$ (115)
where, as before, $\varrho^{(i)},\bar{\varrho}^{(i)}$ are the corresponding density matrices for neutrinos and antineutrinos with momentum $p_{i}$, and $\mathrm{Tr}(\cdots)$ means the trace of the term in front of it.
Finally, the collision terms for the antineutrinos ¯¯$\bar{\mathcal{I}}\equiv\bar{\mathcal{I}}_{\rm sc}+\bar{\mathcal{I}}_{\rm ann}+\bar{\mathcal{I}}_{\nu\nu}$ can be obtained simply by the transformations $\varrho\leftrightarrow\bar{\varrho}$ and $L\leftrightarrow R$.
A.2 Reduced expressions for the comoving collision terms
The comoving collision terms $\tilde{\mathcal{I}}$ in (3.1) are just the above collision terms expressed with the comoving momentum and comoving energies differing by a total factor. In addition, for computational convenience, the nine-dimensional integrals in the collision terms are usually simplified to two-dimensional integrals, using the homogeneous and isotropic condition and the special form of the scattering amplitudes Hannestad:1995rs ; Dolgov:1997mb ; Grohs:2015tfy ; Blaschke:2016xxt . As a result, the collision term can be simplified as
$\tilde{\mathcal{I}}$ $=$ $\frac{G_{F}^{2}}{(2\pi)^{3}y^{2}}\left(\tilde{\mathcal{I}}_{\rm sc}+\tilde{\mathcal{I}}_{\rm ann}+\tilde{\mathcal{I}}_{\nu\nu}\right)\,,$ (116)
~$\tilde{\mathcal{I}}_{\rm sc}$ $=$ $\int{\rm d}y_{2}{\rm d}y_{3}\frac{y_{2}}{E_{2}}$
$\left\{\left(\Pi_{2}^{s}(y,y_{4})+\Pi_{2}^{s}(y,y_{2})\right)\left[F_{\mathrm{sc}}^{LL}(\nu^{(1)},e^{(2)},\nu^{(3)},e^{(4)})+F_{\mathrm{sc}}^{RR}(\nu^{(1)},\bar{e}^{(2)},\nu^{(3)},\bar{e}^{(4)})\right]\right.$
$\left.-2(x^{2}+\delta m_{e}^{2})\Pi_{1}^{s}(y,y_{3})\left[F_{\mathrm{sc}}^{LR}(\nu^{(1)},e^{(2)},\nu^{(3)},e^{(4)})+F_{\mathrm{sc}}^{LR}(\nu^{(1)},\bar{e}^{(2)},\nu^{(3)},\bar{e}^{(4)})\right]\right\}\,,$
~$\tilde{\mathcal{I}}_{\rm ann}$ $=$ $\int{\rm d}y_{2}{\rm d}y_{3}\frac{y_{3}}{E_{3}}$
$\left\{\Pi_{2}^{a}(y,y_{4})F_{\rm ann}^{LL}(\nu^{(1)},\bar{\nu}^{(2)},e^{(3)},\bar{e}^{(4)})+\Pi_{2}^{a}(y,y_{3})F_{\rm ann}^{RR}(\nu^{(1)},\bar{\nu}^{(2)},e^{(3)},\bar{e}^{(4)})\right.$
$\left.+(x^{2}+\delta m_{e}^{2})\Pi_{1}^{a}(y,y_{2})\left[F_{\rm ann}^{RL}(\nu^{(1)},\bar{\nu}^{(2)},e^{(3)},\bar{e}^{(4)})+F_{\rm ann}^{LR}(\nu^{(1)},\bar{\nu}^{(2)},e^{(3)},\bar{e}^{(4)})\right]\right\}\,,$
~$\tilde{\mathcal{I}}_{\nu\nu}$ $=$ $\frac{1}{4}\int{\rm d}y_{2}{\rm d}y_{3}\,$
$\left\{\Pi_{2}^{\nu}(y,y_{2})F_{\nu\nu}(\nu^{(1)},\bar{\nu}^{(2)},\nu^{(3)},\bar{\nu}^{(4)})+\Pi_{2}^{\nu}(y,y_{4})F_{\nu\bar{\nu}}(\nu^{(1)},\bar{\nu}^{(2)},\nu^{(3)},\bar{\nu}^{(4)})\right\}\,,$
where $E^{2}_{i}=\sqrt{x^{2}+y_{i}^{2}+\delta m_{e}^{2}}$ is the comoving energy for electrons/positrons taking into account the mass contribution induced by FTQED corrections (see Appendix B for more details) and the collision kernels in the collision terms are
$\Pi_{1}^{s}(y,y_{3})$ $=$ $y\,y_{3}\,D_{1}+D_{2}(y,y_{3},y_{2},y_{4}),$ (120)
$\Pi_{1}^{a}(y,y_{2})$ $=$ $y\,y_{2}\,D_{1}-D_{2}(y,y_{2},y_{3},y_{4}),$ (121)
$\Pi_{2}^{s}(y,y_{2})/2$ $=$ $y\,E_{2}\,y_{3}\,E_{4}\,D_{1}+D_{3}-y\,E_{2}D_{2}(y_{3},y_{4},y,y_{2})$
$-y_{3}\,E_{4}D_{2}(y,y_{2},y_{3},y_{4}),$
$\Pi_{2}^{s}(y,y_{4})/2$ $=$ $y\,E_{2}\,y_{3}\,E_{4}\,D_{1}+D_{3}+E_{2}\,y_{3}D_{2}(y,y_{4},y_{2},y_{3})$
$+y\,E_{4}D_{2}(y_{2},y_{3},y,y_{4}),$
$\Pi_{2}^{a}(y,y_{3})/2$ $=$ $y\,y_{2}\,E_{3}\,E_{4}\,D_{1}+D_{3}+y\,E_{3}D_{2}(y_{2},y_{4},y,y_{3})$
$+y_{2}\,E_{4}D_{2}(y,y_{3},y_{2},y_{4}),$
$\Pi_{2}^{a}(y,y_{4})/2$ $=$ $y\,y_{2}\,E_{3}\,E_{4}\,D_{1}+D_{3}+y_{2}\,E_{3}D_{2}(y,y_{4},y_{2},y_{3})$
$+y\,E_{4}D_{2}(y_{2},y_{3},y,y_{4}),$
$\Pi_{2}^{\nu}(y,y_{2})/2$ $=$ $y\,y_{2}\,y_{3}\,y_{4}\,D_{1}+D_{3}-y\,y_{2}D_{2}(y_{3},y_{4},y,y_{2})-y_{3}\,y_{4}D_{2}(y,y_{2},y_{3},y_{4}),$ (126)
$\Pi_{2}^{\nu}(y,y_{4})/2$ $=$ $y\,y_{2}\,y_{3}\,y_{4}\,D_{1}+D_{3}+y_{2}\,y_{3}D_{2}(y,y_{4},y_{2},y_{3})+y\,y_{4}D_{2}(y_{2},y_{3},y,y_{4}),$ (127)
where the functions $D_{i}$ are defined as Dolgov:1997mb :
$D_{1}(a,b,c,d)$ $=$ $\frac{16}{\pi}\int_{0}^{\infty}\frac{{\rm d}\lambda}{\lambda^{2}}\prod_{i=a,b,c,d}\sin(\lambda i)\,,$ (128)
$D_{2}(a,b,c,d)$ $=$ $-\frac{16}{\pi}\int_{0}^{\infty}\frac{{\rm d}\lambda}{\lambda^{4}}\prod_{i=a,b}\Big{[}\lambda i\cos(\lambda i)-\sin(\lambda i)\Big{]}\prod_{j=c,d}\sin(\lambda j)\,,$ (129)
$D_{3}(a,b,c,d)$ $=$ $\frac{16}{\pi}\int_{0}^{\infty}\frac{\mathrm{d}\lambda}{\lambda^{6}}\prod_{i=a,b,c,d}\Big{[}\lambda i\cos(\lambda i)-\sin(\lambda i)\Big{]}\,.$ (130)
Again, the comoving collision terms for the antineutrinos ¯~¯~$\tilde{\bar{\mathcal{I}}}\equiv\tilde{\bar{\mathcal{I}}}_{\rm sc}+\tilde{\bar{\mathcal{I}}}_{\rm ann}+\tilde{\bar{\mathcal{I}}}_{\nu\nu}$ can be obtained simply by the transformations $\varrho\leftrightarrow\bar{\varrho}$ and $L\leftrightarrow R$.
Appendix B Finite Temperature QED Corrections
The finite temperature QED (FTQED) corrections on neutrino decoupling must be taking into account to obtain accurate density matrices and the resulting $N_{\rm eff}$ Fornengo:1997wa ; Mangano:2001iu ; Bennett:2019ewm . In this work we consider the following three aspects of the FTQED corrections.
B.1 FTQED correction to the continuity equation
The continuity equation of the Universe with the FTQED corrections is usually expressed as a differential equation for the dimensionless photon temperature $z$ and are solved combing with the QKEs in practice,
$\frac{\mathrm{d}z}{\mathrm{d}x}=\cfrac{rJ_{2}(r)+G_{1}(r)-\cfrac{1}{4\pi^{2}z^{3}}{\displaystyle\int_{0}^{\infty}\mathrm{d}y\,y^{3}\sum_{\alpha}\left(\cfrac{\mathrm{d}\varrho_{\alpha\alpha}}{\mathrm{d}x}+\cfrac{\mathrm{d}\bar{\varrho}_{\alpha\alpha}}{\mathrm{d}x}\right)}}{\Big{[}r^{2}J_{2}(r)+J_{4}(r)\Big{]}+G_{2}(r)+\cfrac{2\pi^{2}}{15}}\,,$ (131)
where $r=x/z$, $\frac{\mathrm{d}\varrho_{\alpha\alpha}}{\mathrm{d}x}$ and $\frac{\mathrm{d}\bar{\varrho}_{\alpha\alpha}}{\mathrm{d}x}$ are the diagonal components of the QKEs for neutrinos and antineutrinos. The remaining functions appearing in the equation are defined as
$J_{a}(r)$ $=$ $\frac{1}{\pi^{2}}\int_{0}^{\infty}{\rm d}u\,u^{a}\frac{\exp(\sqrt{u^{2}+r^{2}})}{\Big{[}\exp(\sqrt{u^{2}+r^{2}})+1\Big{]}^{2}}\,,$ (132)
$K_{a}(r)$ $=$ $\frac{1}{\pi^{2}}\int_{0}^{\infty}{\rm d}u\,\frac{u^{a}}{\sqrt{u^{2}+r^{2}}}\,\frac{1}{\exp(\sqrt{u^{2}+r^{2}})+1}\,.$ (133)
and $G_{1,2}$ functions are expanded as a series of powers of the electron charge $e^{2}=4\pi\alpha$ with $\alpha$ the fine structure constant,
$G_{1,2}(x,z)$ $=$ $G_{1,2}^{(2)}(x/z)+G_{1,2}^{(2+\ln)}(x,z)+G_{1,2}^{(3)}(x/z)+\ldots\,.$ (134)
At the $\mathcal{O}(e^{2})$ order we have
$G_{1}^{(2)}(r)$ $=$ $2\pi\alpha\left[\frac{1}{r}\left(\frac{K_{2}}{3}+2K_{2}^{2}-\frac{J_{2}}{6}-K_{2}J_{2}\right)+G_{a}\right]\,,$ (135)
$G_{2}^{(2)}(r)$ $=$ $-8\pi\alpha\left(\frac{K_{2}}{6}+\frac{J_{2}}{6}-\frac{1}{2}K_{2}^{2}+K_{2}J_{2}\right)+2\pi\alpha rG_{a}\,,$ (136)
$G_{a}(r)$ $=$ $\frac{K_{2}^{\prime}}{6}-K_{2}K_{2}^{\prime}+\frac{J_{2}^{\prime}}{6}+K_{2}^{\prime}J_{2}+K_{2}J_{2}^{\prime}\,,$ (137)
where the prime denotes derivative with respect to $r$ and the explicit dependence on $r$ for the $G$ functions are dropped for simplicity. An additional log term at the $\mathcal{O}(e^{2})$ order is
$G_{1}^{(2+\ln)}(x,z)$ $=$ $\frac{e^{2}x}{16\pi^{4}z^{3}}\int\!\!\!\int_{0}^{\infty}{\rm d}y\,{\rm d}k\,\frac{y\,k}{E_{y}E_{k}}\ln\left|\frac{y+k}{y-k}\right|\Bigg{\{}-x\bigg{[}z\Big{(}\partial_{x}\mathcal{N}_{y}\partial_{z}\mathcal{N}_{k}+\mathcal{N}_{y}\partial_{x}\partial_{z}\mathcal{N}_{k}\Big{)}-\mathcal{N}_{y}\partial_{x}\mathcal{N}_{k}\bigg{]}$ (138)
$\qquad\qquad\qquad\qquad-\mathcal{N}_{y}\mathcal{N}_{k}-z\mathcal{N}_{y}\partial_{z}\mathcal{N}_{k}+\frac{x^{2}(E_{y}^{2}+E_{k}^{2})}{2E_{y}^{2}E_{k}^{2}}\Big{(}2z\mathcal{N}_{y}\partial_{z}\mathcal{N}_{k}-\mathcal{N}_{y}\mathcal{N}_{k}\Big{)}\Bigg{\}}\,,$
$G_{2}^{(2+\ln)}(x,z)$ $=$ $\frac{e^{2}x^{2}}{16\pi^{4}z^{2}}\int\!\!\!\int_{0}^{\infty}{\rm d}y\,{\rm d}k\,\frac{y\,k}{E_{y}E_{k}}\ln\left|\frac{y+k}{y-k}\right|\partial_{z}\Big{(}\mathcal{N}_{y}\partial_{z}\mathcal{N}_{k}\Big{)}\,,$ (139)
where
$\mathcal{N}_{p}$ $=$ $\frac{2}{e^{E_{p}/z}+1}\,,$ (140)
$\partial_{x}\mathcal{N}_{p}$ $=$ $-\frac{x\,e^{E_{p}/z}\,\mathcal{N}_{p}^{2}}{2z\,E_{p}}\,,$ (141)
$\partial_{z}\mathcal{N}_{p}$ $=$ $\frac{e^{E_{p}/z}\,E_{p}\,\mathcal{N}_{p}^{2}}{2z^{2}}\,,$ (142)
$\partial_{x}\partial_{z}\mathcal{N}_{p}$ $=$ $\frac{x\,e^{E_{p}/z}\,\mathcal{N}_{p}^{2}}{2z^{3}}\left(1-e^{E_{p}/z}\,\mathcal{N}_{p}+\frac{z}{E_{p}}\right)\,.$ (143)
At the $\mathcal{O}(e^{3})$ order we have
$G_{1}^{(3)}(r)$ $=$ $\frac{e^{3}}{4\pi}G_{b}\left\{\frac{1}{r}\Big{(}2J_{2}-4K_{2}\Big{)}-2J_{2}^{\prime}-r^{2}J_{0}^{\prime}-r\Big{(}2K_{0}+J_{0}\Big{)}\right.$ (144)
$\left.-G_{c}\Big{[}r(K_{0}-J_{0})+K_{2}^{\prime}\Big{]}\right\}\,,$
$G_{2}^{(3)}(r)$ $=$ $\frac{e^{3}}{4\pi}G_{b}\left[G_{c}\Big{(}2J_{2}+r^{2}J_{0}\Big{)}-\frac{2}{r}J_{4}^{\prime}-r\Big{(}3J_{2}^{\prime}+r^{2}J_{0}^{\prime}\Big{)}\right]\,,$ (145)
with
$G_{b}(r)$ $=$ $\sqrt{K_{2}+r^{2}\frac{K_{0}}{2}}\,,$ (146)
$G_{c}(r)$ $=$ $\frac{2J_{2}+r^{2}J_{0}}{2\Big{(}2K_{2}+r^{2}K_{0}\Big{)}}\,.$ (147)
B.2 FTQED corrections to the total energies and the pressures
The total pressure and energy density of the universe are also modified by the FTQED corrections:
$P_{\rm tot}$ $=$ $\sum_{i=\gamma,{\nu_{i}},e}P_{i}+\delta P(x,z)\,,$ (148)
$\rho_{\rm tot}$ $=$ $\sum_{i=\gamma,{\nu_{i}},e}\rho_{i}+\delta\rho(x,z)\,,$ (149)
where $\delta P$ and $\delta\rho$ can also be expanded as a series of powers of the electron charge, so that we have Bennett:2019ewm
$\delta P(x,z)$ $=$ $\delta P^{(2)}(x/z)+\delta P^{(2+\ln)}(x,z)+\delta P^{(3)}(x/z)+\ldots\,,$ (150)
$\delta P^{(2)}(r)$ $=$ $-e^{2}z^{4}\,K_{2}\left(\frac{1}{6}+\frac{K_{2}}{2}\right)\,,$ (151)
$\delta P^{(2+\ln)}(x,z)$ $=$ $\frac{e^{2}x^{2}}{16\pi^{4}}\int\!\!\!\int_{0}^{\infty}{\rm d}y\,{\rm d}k\,\frac{y\,k}{E_{y}E_{k}}\ln\left|\frac{y+k}{y-k}\right|\mathcal{N}_{y}\,\mathcal{N}_{k}\,,$ (152)
$\delta P^{(3)}(r)$ $=$ $\frac{2e^{3}z^{4}}{3\pi}\left(K_{2}+\frac{r^{2}}{2}K_{0}\right)^{3/2}\,,$ (153)
for the pressure, while
$\delta\rho(x,z)$ $=$ $\delta\rho^{(2)}(x/z)+\delta\rho^{(2+\ln)}(x,z)+\delta\rho^{(3)}(x/z)+\ldots\,,$ (154)
$\delta\rho^{(2)}(r)$ $=$ $e^{2}z^{4}\left(\frac{K_{2}^{2}}{2}-\frac{K_{2}+J_{2}}{6}-K_{2}J_{2}\right)\,,$ (155)
$\delta\rho^{(2+\ln)}(x,z)$ $=$ $\frac{e^{2}x^{2}}{16\pi^{4}}\int\!\!\!\int_{0}^{\infty}{\rm d}y\,{\rm d}k\,\frac{y\,k}{E_{y}E_{k}}\ln\left|\frac{y+k}{y-k}\right|\mathcal{N}_{y}\Big{(}2z\partial_{z}\mathcal{N}_{k}-\mathcal{N}_{k}\Big{)}\,,$ (156)
$\delta\rho^{(3)}(r)$ $=$ $\frac{e^{3}z^{4}}{\pi}\left(K_{2}+\frac{r^{2}}{2}K_{0}\right)^{1/2}\left(J_{2}+\frac{r^{2}}{2}J_{0}\right)\,.$ (157)
for the energy densities.
B.3 FTQED correction to the electron mass
In addition, the mass of electron/positrons are also modified by the FTQED correction, which is mainly used in the collision terms. In the comoving coordinates, the additional FTQED contribution to the electron mass can be written as Fornengo:1997wa ; Mangano:2001iu ; Bennett:2019ewm :
$\delta m_{e}^{2}(x,y,z)=\frac{2\pi\alpha z^{2}}{3}+\frac{4\alpha}{\pi}\int_{0}^{\infty}{\rm d}k\,\frac{k^{2}}{E_{k}}\frac{1}{e^{E_{k}/z}+1}-\frac{x^{2}\alpha}{\pi y}\int_{0}^{\infty}{\rm d}k\,\frac{k}{E_{k}}\log\left|\frac{y+k}{y-k}\right|\mathcal{N}_{k}\,,$ (158)
so that the comoving electron mass in the collision terms is replaced using $x^{2}\rightarrow x^{2}+\delta m_{e}^{2}$. Finally, we note that the the log term that depends on $y$ in the FTQED corrections to the continuity equation, the total energy densities and the pressure, as long as the electron mass are ignored in the computation, as it is shown that their contribution is negligible Froustey:2020mcq ; Bennett:2020zkv .